### v1

# Integrating Large Language Models into Recommendation via Mutual Augmentation and Adaptive Aggregation

Abstract.
Conventional recommendation methods have achieved notable advancements by harnessing collaborative or sequential information from user behavior. Recently, large language models (LLMs) have gained prominence for their capabilities in understanding and reasoning over textual semantics, and have found utility in various domains, including recommendation. Conventional recommendation methods and LLMs each have their own strengths and weaknesses. While conventional methods excel at mining collaborative information and modeling sequential behavior, they struggle with data sparsity and the long-tail problem. LLMs, on the other hand, are proficient at utilizing rich textual contexts but face challenges in mining collaborative or sequential information. Despite their individual successes, there is a significant gap in leveraging their combined potential to enhance recommendation performance.
In this paper, we introduce a general and model-agnostic framework known as Large Language model with mutual augmentation and adaptive aggregation for Recommendation (Llama4Rec). Llama4Rec synergistically integrates conventional and LLM-based recommendation models. Llama4Rec proposes data augmentation and prompt augmentation strategies tailored to enhance the conventional model and the LLM respectively. An adaptive aggregation module is adopted to combine the predictions of both kinds of models to refine the final recommendation results. Empirical studies on three real-world datasets validate the superiority of Llama4Rec, demonstrating its consistent and significant improvements in recommendation performance over baseline methods.
${}^{\dagger}$oresponding Author
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: The 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14-18, 2024; Washington D.C., USA
1. Introduction
Recommender systems have emerged as crucial solutions for mitigating the challenge of information overload (He et al., 2023a, b; Luo et al., 2023a, 2022a). Recommender systems encompass a multitude of tasks, such as rating prediction (Steck, 2013; Khan et al., 2021) and top-$k$ recommendation (Le and Lauw, 2021; Luo et al., 2023b). The top-$k$ recommendation, which encompasses collaborative filtering-based direct recommendation (He et al., 2017, 2020), sequential recommendation (Kang and McAuley, 2018; Sun et al., 2019; Chen et al., 2023), and more, has found wide applications in various areas. However, recommender systems still suffer from the data sparsity and long-tail problem. Data sparsity arises from sparse user-item interactions, making the task of accurately capturing user preferences more challenging. The long-tail problem further intensifies data sparsity issue, as a substantial number of less popular items (i.e., long-tail items) are infrequently interacted with, leading to inadequate data for effective model training and compromised recommendation quality.
In recent years, Large Language Models (LLMs) have emerged, exhibiting exceptional capabilities in language understanding, text generation, and complex reasoning tasks (OpenAI, 2023; Touvron et al., 2023a, b; Wang et al., 2023; Zhou et al., 2023). Recent studies have started exploring their applicability in recommender systems (Liu et al., 2023; Bao et al., 2023; Zhang et al., 2023c). For example, Liu et al. employed ChatGPT with in-context learning (ICL) for various recommendation tasks (Liu et al., 2023). Further progress has been achieved by adopting the instruction tuning technique (Ouyang et al., 2022; Longpre et al., 2023) to align general-purpose LLMs with recommendation tasks for improved performance (Zhang et al., 2023c; Bao et al., 2023). For instance, TALLRec (Bao et al., 2023) reformulates the recommendation problem as a binary classification task and introduces an effective instruction fine-tuning framework for adapting the LLaMA model (Touvron et al., 2023b). However, these LLM-based recommendation methods may not perform optimally as they do not harness the collaborative or sequential information captured by conventional recommendation models.
Conventional recommendation models and LLM-based recommendation methods each have their respective strengths and weaknesses. Conventional methods excel in mining collaborative information and modeling sequential behaviors, while LLMs are proficient in leveraging rich textual contexts. As such, the integration of LLMs into recommender systems presents a significant opportunity to amalgamate the advantages of both methodologies while circumventing their respective shortcomings. There have been initial attempts to harness the strengths of both conventional and LLM-based recommenders (Geng et al., 2022; Zhang et al., 2023b; Zheng et al., 2023; Wei et al., 2023). Some efforts have sought to integrate collaborative/sequential information by enabling LLMs to comprehend user/item ID information (Geng et al., 2022; Zhang et al., 2023b; Zheng et al., 2023). For instance, a concurrent study by Zhang et al. encoded the semantic embedding into the prompt (Zhang et al., 2023b) and send it to LLM. On the other hand, some research works have aimed to augment conventional models using LLMs via data or knowledge augmentation (Wei et al., 2023; Xi et al., 2023). LLMRec (Wei et al., 2023) enhances recommender systems by deploying LLMs to augment the interaction graph, thereby addressing the challenges of data sparsity and low-quality side information.
However, existing methods have several limitations. Firstly, current methods lack generalizability. The strategy of integrating ID information proves challenging to generalize across different domains and necessitates additional training. The current data augmentation method is not universally applicable, as it only addresses a limited number of recommendation scenarios. Secondly, current research primarily focuses on the integration at the data-level (e.g., data augmentation) or model-level (e.g., make LLM understand ID semantics), leaving the result-level integration largely unexplored. Lastly, there is an absence of a comprehensive framework that combines and integrates these methods into a single construct. In light of these limitations, our objective is to explore the integration of conventional recommendation models and LLM-based recommendation methods in depth to address the above limitations and enhance recommendation performance.
In this paper, we introduce a general framework known as Large language model with mutual augmentation and adaptive aggregation for Recommendation, referred to as Llama4Rec, for brevity. The core idea of Llama4Rec is to allow conventional recommendation models and LLM-based recommendation models to mutually augment each other, followed by an adaptive aggregation of the augmented models to yield more optimized results. Specifically, Llama4Rec performs data augmentation for conventional recommendation models by leveraging instruction-tuned LLM to alleviate the data sparsity and long-tail problem. The data augmentation is tailored with different strategies depending on the recommendation scenarios. Furthermore, we use conventional recommendation models to perform prompt augmentation for LLMs. The prompt augmentation includes enriching collaborative information from similar users and providing prior knowledge from the conventional recommendation model within the prompt. We also propose an adaptive aggregation module that merges the predictions of the LLM and conventional models in an adaptive manner. This module is designed as a simple yet effective way to combine the strengths of both models and refine the final recommendation results. We conduct empirical studies on three real-world datasets, encompassing three different recommendation tasks, to validate the superiority of our proposed method. The results consistently demonstrate its superior performance over baseline methods, highlighting notable improvements in recommendation performance.
In a nutshell, the contributions of this work are threefold.
  • We introduce Llama4Rec, a general and model-agnostic framework to integrate LLM into conventional recommendation models. Llama4Rec performs the data augmentation for conventional models to alleviate the data sparsity problem and improve model performance. The prompt augmentation is applied to LLM for leveraging the information captured by the conventional models.
  • Llama4Rec employs an adaptive aggregation approach to combine the prediction from the conventional recommendation model and LLM for improved recommendation performance via leveraging and merging the information by both kinds of models.
  • To validate the effectiveness of Llama4Rec, we conduct extensive experiments on three real-world datasets across three diverse recommendation tasks. The empirical results demonstrate that Llama4Rec outperforms existing baselines, exhibiting notable improvements across multiple performance metrics.
2. Related Work
2.1. Conventional Recommendation Methods
Conventional recommendation methods serve as the cornerstone for the contemporary landscape of recommender systems (Zhang et al., 2019). Representative recommendation tasks include rating prediction, collaborative filtering-based direct recommendation, and sequential recommendation, where the latter two are usually formulated as top-$k$ recommendation problems. Specifically, one of the seminal techniques is the use of matrix factorization for rating prediction, popularized by methods such as Singular Value Decomposition (SVD) (Koren et al., 2009). Collaborative filtering (CF) is another commonly used technique for the recommender systems (Adomavicius and Tuzhilin, 2005). Recent advancements have evolved CF techniques into more complex neural network architectures and graph-based models (He et al., 2017; Wang et al., 2019) to enhance the model performance. Sequential recommendation models incorporate temporal patterns into the recommendation pipeline. Techniques such as recurrent neural networks have been adapted for this purpose (Hidasi et al., 2015). Recent research focuses on applying attention mechanisms to further refine these models, leading to a noteworthy boost in performance (Sun et al., 2019; Kang and McAuley, 2018).
Although conventional recommendation techniques are well-suited for capturing latent information associated with users and items, they often require a substantial amount of user-item interactions to provide accurate recommendations, which limits their effectiveness in data sparse and long-tail scenarios (Park and Tuzhilin, 2008).
Figure 1. (i) The overall framework architecture of the proposed Llama4Rec consists of two main components: mutual augmentation and adaptive aggregation. The mutual augmentation includes data augmentation and prompt augmentation. (ii) Illustration of the data augmentation process encompasses three diverse recommendation scenarios. (iii) The pipeline of the adaptive aggregation module, which merges the predictions from the conventional recommendation model and the LLM.
2.2. Large Language Model for Recommendation
LLMs have brought a paradigm shift in numerous areas of machine learning, including recommendation methods (Fan et al., 2023). One of the most compelling advantages of LLM-based recommendation methods is their capacity for contextual understanding and in context learning (Dong et al., 2022). Inspired by this, reference (Liu et al., 2023) utilized ChatGPT across diverse recommendation tasks and found it to be effective in specific contexts, underpinned by robust experiments and human evaluations. Similarly, Wang et al. (Wang and Lim, 2023) introduced a novel zero-shot technique for next-item recommendations, further substantiating the utility of LLMs in this arena. However, it is noteworthy that these methods do not consistently demonstrate a marked improvement over conventional recommendation algorithms, which is largely attributable to the inherent misalignment between the general-purpose capabilities of LLMs and the specialized requirements of the recommendation task. To address this issue, recent studies further attempt to instruct tuning the specific LLM to align with human preference (Bao et al., 2023; Zhang et al., 2023c). Typically, these approaches involve creating an instruction tuning dataset in line with recommendation tasks, which is then used to tune the LLM for recommendation. Such methodologies have demonstrated improved performance in generating more aligned and accurate recommendations.
Nevertheless, while LLMs excel at capturing intricate textual patterns, they may encounter challenges in comprehensively encoding user and item collaborative or sequential information. Though some concurrent studies (Zhang et al., 2023b; Zheng et al., 2023) aim to address this gap, they often lack in terms of generalizability and comprehensibility. In response to this challenge, we propose a novel framework designed to mitigate this issue.
3. Preliminary
We consider a recommender system with a set of users, denoted $\mathcal{U}=\{u_{1},u_{2},\ldots,u_{n}\}$, and a set of items, denoted =$\mathcal{I}=\{i_{1},i_{2},\ldots,i_{m}\}$. The rating prediction task aims to estimate the unknown values of $r_{ui}$ in the user-item interaction matrix $R\in\mathbb{R}^{n\times m}$, where each entry $r_{ui}$ is the rating assigned by user $u$ to item $i$. Different from rating prediction, top-$k$ recommendation focuses on identifying a subset of items $\mathcal{S}_{u}\subset\mathcal{I}$ for each user $u$. The subset is chosen to maximize a user-specific utility $U(u,\mathcal{S})$ with the constraint $|\mathcal{S}|=k$, which can be formally expressed as:
(1) $\mathcal{S}_{u}={\arg\max}_{\mathcal{S}\subset\mathcal{I},|\mathcal{S}|=k}U(u,\mathcal{S}).$
In the context of LLM-based recommendation methods, let ${LLM}$ represent the original LLM. These kinds of methods first utilize prompt to interpret the recommendation task into natural language. The LLM-based recommendation for user $u$ with in-context learning is denoted by ${r}_{u}={LLM}(\mathcal{P}_{u})$ where $\mathcal{P}_{u}$ is the recommendation prompt for user $u$. The recommendation prompt could either ask LLM to predict a rating towards a target item, or rank candidate items derived by retrieval models in top-k recommendation. To instruction fine-tune LLM, a dedicated dataset $\mathcal{D}_{ins}$ consisting of various instructions is utilized. The resulting instruction-tuned LLM is denoted as ${LLM}^{\prime}$. Therefore, the recommendation process in the fine-tuned model can be succinctly represented as $r_{u}={LLM}^{\prime}(\mathcal{P}_{u})$.
4. Methodology
4.1. Overview
Figure 1 depicts the architecture of Llama4Rec, which consists of three components: data augmentation, prompt augmentation, and adaptive aggregation. More specifically, Llama4Rec leverages an instruction-tuned LLM to enhance conventional recommendation systems through data augmentation. The specific data augmentation strategies for different recommendation situations are detailed in Section 4.2. In addition, we employ conventional recommendation models to augment the LLM via prompt augmentation, with details in Section 4.3. To further refine the predictions of the conventional model and the LLM, we propose a simple yet effective adaptive aggregation module in Section 4.4. Lastly, we describe the training strategy for LLM in Section 4.5.
4.2. Data Augmentation for Conventional Recommendation Model
We design ad-hoc data augmentation strategies for different recommendation scenarios to mitigate prevalent issues of data sparsity and the long-tail problem. This design is motivated by the fact that data distribution and tasks significantly vary across different recommendation scenarios. In the context of direct recommendation, we capitalize on the power of the instruction-tuned LLM to predict items that a user may like or dislike. We form pairs of these items to calculate the Bayesian Personalized Ranking (BPR) (Rendle et al., 2012) loss. For sequential recommendation, we harness the capabilities of the instruction-tuned LLM to predict items that are highly preferred by the user. These predicted items are then randomly inserted into the sequence of items the user has interacted with. For rating prediction, we utilize the LLM to extract valuable side information (i.e., missing attributes), which is then seamlessly integrated as additional features within the training data.
4.2.1. Data Augmentation for Direct Recommendation
For direct recommendation, the Bayesian Personalized Ranking (BPR) loss is commonly used to optimize the model (Rendle et al., 2012). The objective of BPR is to maximize the score difference between correctly recommended items and incorrectly recommended items, thereby improving the accuracy of recommendations. The BPR loss is defined as:
(2) $\mathcal{L}_{BPR}=-\sum_{(u,i,j)\in\mathcal{D}}\log\sigma(\hat{y}_{ui}-\hat{y}_{uj}),$
where $(u,i,j)$ refers to a triple of user-item pairs, and the user $u$ has interacted with item $i$ (positive item) and item $j$ (negative item). $\mathcal{D}$ represents the set of such user-item pairs in the training data. $\hat{y}_{ui}$ denotes the predicted score or preference of user $u$ for item $i$.
[…]
(3) $\mathcal{L}_{BPR}^{\prime}=-\sum_{(u,i,j)\in\mathcal{D}^{\prime}}\log\sigma(\hat{y}_{ui}-\hat{y}_{uj}).$
This data augmentation strategy leverages the power of the instruction tuned LLM to enhance the performance of the direct recommendation model.
4.2.2. Data Augmentation for Sequential Recommendation
[…]
By including additional items predicted by the LLM, we can enrich the sequence of items for each user, providing a more comprehensive representation of the user’s preferences. This, in turn, can enhance the performance of the conventional recommendation model, leading to more accurate recommendations.
4.2.3. Data Augmentation for Rating Prediction
In rating prediction tasks, we introduce the use of in-context learning (ICL) in LLMs to provide side information. This is primarily due to the fact that recommendation datasets may contain incomplete information. For instance, the popular Movielens dataset (Harper and Konstan, 2015) lacks information about the director of the movies, which can hinder the performance of a conventional rating prediction model. To mitigate this issue, we leverage the extensive world knowledge contained in an LLM. We prompt the LLM to provide side information, acting as additional attributes for users/items.
[…]
Top-k Recommendation Prompt Example:
Instruction: Rank the candidate movies based on user historical interactions and make the top k recommendations.
Interaction History: Beyond Rangoon (1995); Alien (1979); Hollow Reed (1996); Primary Colors (1998); …; Birds, The (1963)
Candidate Items: Last Dance (1996); Remains of the Day, The (1993); Assassins (1995); …; Fatal Instinct (1993)
Similar User Interaction History: L.A. Confidential (1997); Apt Pupil (1998); Kolya (1996); …; Star Wars (1977)
Conventional Model Prediction: Remains of the Day, The (1993); Addiction, The (1995); …; Fugitive, The (1993)
Output: Fugitive, The (1993); Angel Baby (1995); …; Remains of the Day, The (1993)
Rating Prediction Prompt Example:
Instruction: Predict the rating of a target movie based on the user’s historical movie ratings.
Rating History: Independence Day (1996): 3; Grosse Fatigue (1994): 3; Face/Off (1997): 4; …; Shall We Dance? (1996): 3
Candidate Item: Pink Floyd - The Wall (1982)
Similar User Rating History: L.A. Confidential (1997): 3; Apt Pupil (1998): 4; …; English Patient, The (1996): 3
Conventional Model Prediction: 3.2
Output: 3
Figure 2. Examples of instructions for top-k recommendation and rating prediction. The prompt augmentation component is underlined. To improve readability and facilitate better experimental evaluation, we introduce certain modifications to the original instructions employed in our experiments.
4.3. Prompt Augmentation for Large Language Model
Previous works (Bao et al., 2023; Zhang et al., 2023c) instruction tuning LLM for recommendation in a standard manner. However, these methods can be sub-optimal due to the challenges of distinguishing users based solely on text-based prompt descriptions. Although some concurrent studies (Zhang et al., 2023b; Zheng et al., 2023) incorporate unique identifiers to differentiate users, these approaches require complex semantic understanding of IDs and additional training, limiting their generalizability.
In this section, we introduce two text-based prompt augmentation strategies for LLM-based recommendations, i.e., we incorporate additional information within the prompt to enhance the model performance. First, we propose prompt augmentation with similar users, identifying users with analogous preferences to enrich the prompt, thereby enhancing the LLM’s ability to leverage collaborative information and generate personalized recommendations. Second, we propose prompt augmentation with conventional model prediction, providing prior knowledge to guide the LLM toward recommendations that align with user preferences. Collectively, these strategies harness the strengths of both LLMs and conventional recommendation models, ensuring generalizability across a wide range of recommendation scenarios. The illustration of prompt augmentation is underlined in Figure 2.
4.3.1. Prompt Augmentation with Collaborative Information from Similar User
To incorporate collaborative information within the prompt and facilitate LLM reasoning, we introduce a prompt augmentation strategy with similar user. Initially, we utilize a pre-trained conventional recommendation model to acquire embeddings for each user. These embeddings represent users in a latent space, which encapsulates their preferences and behaviors. Specifically, for a user $u$, in conjunction with a conventional recommendation model $\mathcal{M}_{c}$, we use $\mathcal{M}_{c}$ to obtain embeddings for each user, denoted as $\{e_{1},...,e_{n}\}$. These embeddings encapsulate the preferences and behaviors of the users, serving as a compact representation of the users in a latent space. We then calculate the similarity between these embeddings in the latent space. Various measurements, such as cosine similarity, Jaccard similarity, and Euclidean distance, could be employed in this context. In this paper, we calculate the cosine similarity to measure how closely two vectors align, denoted as:
(4) $sim(u,v)=\frac{e_{u}\cdot e_{v}}{||e_{u}||\cdot||e_{v}||},$
where $e_{u}$ and $e_{v}$ are the embeddings of user $u$ and $v$, respectively, and $u,v\in\mathcal{U}$. The |||||$\cdot$||| | ⋅ | | denotes the Euclidean norm and $\cdot$ denotes the dot product. We identify the pair of users $(u,v)$ that have the highest similarity, indicating that they are the most similar in terms of their preferences and behaviors. We then use the items interacted with by the most similar user to enrich the prompt for the target user. This strategy leverages the collaborative information gleaned from similar users to generate more relevant and accurate prompts, thereby enhancing the recommendation performance of the LLM.
4.3.2. Prompt Augmentation with Prior Knowledge from Conventional Recommendation Model Prediction
To enable the LLM to leverage information captured by conventional models, we propose a prompt augmentation method that incorporates information from conventional recommendation models. More specifically, the augmented prompt is formed by concatenating the original prompt with the prediction from the conventional recommendation model in natural language form. It’s important to note that the prediction from the conventional model varies depending on the recommendation scenarios and base models. Through augmenting prompts with predictions from conventional recommendation models, our method integrates collaborative or sequential information captured by these models, thereby enhancing the LLM’s contextual understanding and reasoning capabilities and resulting in better recommendation performance.
Notably, unlike ID-based methods such as (Zhang et al., 2023b; Zheng et al., 2023), our approach relies entirely on text, enabling easy adaptation to new situations. Also, the prompt augmentation could be used as a plug-and-play component for recommendation with closed source LLM, such as the GPT-4 model (OpenAI, 2023).
4.4. Adaptive Aggregation
We endeavor to aggregate the outputs of LLM and conventional recommendation models at the result level for improved performance, considering the disparate model structures. However, indiscriminate aggregation of model predictions can potentially lead to suboptimal results. Conventional recommendation models, known for their susceptibility to the long tail issue, often struggle when dealing with the tail segment. In contrast, LLMs, by leveraging contextual information, are able to maintain a relatively uniform performance across all segments. Motivated by these observations, we first define the long-tail coefficient and subsequently adaptively aggregate the predictions from both model types.
We first define the long-tail coefficient $\ell_{u}$ for user $u$ to quantify where the user is located in the tail of the distribution. The long-tail coefficient is defined as follows:
(5) $\ell_{u}=\text{log}\left(N({u})+1\right),$
where $N({u})$ is the number of interaction for user $u$. A lower long-tail coefficient value indicates that the user has fewer feedback.
While the overarching architecture remains consistent, the implementation details are different for the two tasks considered, namely rating prediction and top-$k$ recommendation.
4.4.1. Adaptive Aggregation for Rating Prediction
For the rating prediction task, we employ an instruction-tuned LLM to predict user-item utility scores directly. This approach incorporates the understanding of complex semantics and context by the LLM, which might be overlooked by traditional models. Similarly, conventional recommendation methods leverage collaborative information and user/item features for predicting the user rating. Specifically, the utility weight for user $u$, denoted as $U_{u}$, is directly set as its user rating. Subsequently, the LLM is engaged to predict the rating, symbolized as $U_{LLM}$. There are various methods to derive a final result based on the utility scores, such as training a neural network to process the utility scores from LLM and conventional models, yielding a final output via learning the complex reflection. However, for the sake of simplicity in this paper, we adopt a simple yet effective linear interpolation approach. The final utility score for a user $u$ amalgamates the values from both models, represented as:
(6) $U_{u}=\alpha_{u}{U}_{LLM}+(1-\alpha_{u})U_{Rec},$
where $\alpha_{u}$ is the adaptive parameter to control the weight for each model’s utility value for user $u$. We define the $\alpha_{u}$ as:
(7) $\alpha_{u}=\text{max}\left(\frac{\ell_{max}-\ell_{u}}{\ell_{max}-\ell_{min}},\alpha_{2}\right)\cdot\alpha_{1},$
where $\ell_{max}$ and $\ell_{min}$ are the maximum and minimum long-tail coefficients of the users, respectively, $\alpha_{1}$ is a hyper-parameter that controls the weight, and $\alpha_{2}<1$ is a cut-off weight. From Equation (7), we can observe that for user $u$, the further they are positioned in the long tail (i.e., the fewer items they have interacted with), the lower is the value of $\ell_{u}$ and the higher is the value of $\alpha_{u}$. As a result, in Equation (6), the weight of the utility score from the LLM model becomes more pronounced. This aligns with the motivation we previously discussed.
4.4.2. Adaptive Aggregation for Top-$k$ Recommendation
For the top-$k$ recommendation task, the LLM is employed to re-rank the item list generated by a conventional recommendation model. Specifically, from conventional recommendation methods, we curate a top-ranked list comprising $k^{\prime}$ items, denoted as $\{i_{1},...,i_{k^{\prime}}\}$. Each item in this list is assigned a utility weight, ${U}^{i}_{Rec}=-s\cdot\mathcal{C}$, where $\mathcal{C}$ is a constant and $s$ represents the position of item $i$, i.e., $s\in\{1,...,k^{\prime}\}$. A higher utility weight indicates a stronger inclination of the user’s preference. For listwise comparison conducted by the LLM, the process begins by using the LLM to directly output the predicted order of these candidate items. Then we assign utility scores for items at each position, denoted as $U_{1},{U}_{2},...,U_{k^{\prime}}$, where $U_{1}\geq U_{2}\geq...\geq U_{k^{\prime}}$. The final utility score for an item amalgamates the values from both the original rating and the LLM prediction, similar to the Equation (6).
4.5. Training Strategy for LLM
4.5.1. Instruction Tuning Dataset Construction
This section details the creation of an instruction-tuning dataset that encompasses two types of recommendation tasks catering to top-$k$ recommendation and rating prediction scenarios. A depiction of these two tasks, specifically referred to as listwise ranking and rating prediction, can be found in Figure 2. It is noteworthy that we also employ the LLM to execute pointwise ranking within top-$k$ recommendation scenarios, i.e., utilizing LLM to predict ratings for each item within the top-k recommendations and sorting the predicted ratings to derive the final result.
4.5.2. Optimization via Instruction Tuning
In this work, we perform full parameter instruction tuning to optimize LLMs using generated instruction data. Due to our need for customization, we chose LLaMA-2 (Touvron et al., 2023b), an open-source, high-performing LLM, which permits task-specific fine-tuning. During supervised fine-tuning, we apply a standard cross-entropy loss following Alpaca (Taori et al., 2023). The training set $\mathcal{D}_{ins}$ consists of instruction input-output pairs $(x,y)$, which have been represented in natural language. The objective is to fine-tune the pre-trained LLM by minimizing the cross-entropy loss, formalized as:
(8) $\min_{\Theta}\sum_{(x,y)\in\mathcal{D}_{ins}}\sum_{t=1}^{|y|}-\log P_{\Theta}\left(y_{t}\mid x,y_{[1:t-1]}\right),$
where $\Theta$ are the original parameters for LLM, $P_{\Theta}$ is the conditional probability, $|y|$ is the number of tokens in $y$, $y_{t}$ is the $t$-th token in the target output $y$, and $y_{[1:t-1]}$ represents tokens preceding $y_{t}$ in $y$. By minimizing this loss function, the model fine-tunes its parameters $\Theta$ to adapt to the specifics of the new instruction tuning dataset $\mathcal{D}_{ins}$, while leveraging the general language understanding and reasoning that has been acquired during pre-training (Zhang et al., 2023a). In this manner, LLM can capture the user’s preferences for items expressed in natural language, facilitating diverse recommendation tasks, including top-k recommendation and rating prediction.
5. Experiment
In this section, we present a thorough empirical evaluation to validate the effectiveness of our proposed framework. Specifically, our objective is to investigate whether the incorporation of our proposed Llama4Rec could enhance existing recommendation models. The overarching goal is to answer the following research questions:
  • RQ1: Does our proposed Llama4Rec framework enhance the performance of existing recommendation models?
  • RQ2: How do the various modules in Llama4Rec affect the recommendation performance?
  • RQ3: How do different hyper-parameters impact the overall performance of the framework?
5.1. Experiment Setup
5.1.1. Dataset
Following (Bao et al., 2023), we rigorously evaluate the performance of our proposed framework by employing three heterogeneous, real-world datasets. MovieLens [^1] https://grouplens.org/datasets/movielens/ (Harper and Konstan, 2015) serve as benchmark datasets in the realm of movie recommendation methods. We employ two variants of the dataset: MovieLens-100K (ML-100K) and MovieLens-1M (ML-1M). The former consists of approximately 100,000 user-item ratings, while the latter scales up to roughly 1,000,000 ratings. BookCrossing [^2] Due to the absence of timestamp data, we synthesize historical interactions through random sampling. (Ziegler et al., 2005) includes user-generated book ratings on a scale of 1 to 10, alongside metadata such as ‘Book-Author’ and ‘Book-Title’. We employ LLM to augment the ‘director’ and ‘star’ features for ML-100K and ML-1M datasets, and augment the ‘genre’ and ‘page length’ features for the BookCrossing dataset. To ensure the data quality, we adopt the 5-core setting, i.e., we filter unpopular users and items with fewer than five interactions for the BookCrossing dataset. The key characteristics of these datasets are delineated in Table 1.
Table 1. Dataset Description.
ML-100K ML-1M BookCrossing # of User 943 6,040 6,851 # of Item 1,682 3,706 9,085 # of Rating 100,000 1,000,209 115,219 Density 0.063046 0.044683 0.001851 User Features Gender, ZipCode, Gender, ZipCode, Location, Age Occupation, Age Occupation, Age Item Features Title, Genres Title, Genres Title, Author, Year Year, Publisher Augmented Movie Director, Movie Director, Book Genres, Features Movie Star Movie Star Page Length
5.1.2. Evaluation Metrics.
Aligning with (He et al., 2020; Sun et al., 2019), for the top-$k$ recommendation task, we turn to two well-established metrics: Hit Ratio (HR) and Normalized Discounted Cumulative Gain (NDCG), denoted by H and N, respectively. In our experiments, $k$ is configured to be either 3 or 5 for a comprehensive evaluation, similar to the experiment setting in (Zhang et al., 2023c). In accordance with (Fan et al., 2019), we employ Root Mean Squared Error (RMSE) and Mean Absolute Error (MAE) as evaluation metrics to ascertain the performance of the rating prediction task.
Table 2. Performance achieved by different direct recommendation methods.
Backbone Method ML-100K ML-1M BookCrossing H@3 $\uparrow$ N@3 $\uparrow$ H@5 $\uparrow$ N@5 $\uparrow$ H@3 $\uparrow$ N@3 $\uparrow$ H@5 $\uparrow$ N@5 $\uparrow$ H@3 $\uparrow$ N@3 $\uparrow$ H@5 $\uparrow$ N@5 $\uparrow$ MF Base 0.0455 0.0325 0.0690 0.0420 0.0255 0.0187 0.0403 0.0248 0.0294 0.0227 0.0394 0.0269 IFT 0.0546 0.0388 0.0790 0.0488 0.0242 0.0175 0.0410 0.0244 0.0247 0.0177 0.0377 0.0230 Llama4Rec 0.0645* 0.0474* 0.0919* 0.0588* 0.0281* 0.0203* 0.0433* 0.0265* 0.0365* 0.0284* 0.0462* 0.0324* Impro. 18.13% 22.16% 16.33% 20.49% 10.20% 8.56% 5.61% 6.85% 24.15% 25.55% 17.26% 20.45% LightGCN Base 0.0492 0.0343 0.0744 0.0447 0.0283 0.0203 0.0432 0.0264 0.0358 0.0272 0.0480 0.0322 IFT 0.0537 0.0381 0.0846 0.0507 0.0268 0.0193 0.0441 0.0263 0.0287 0.0202 0.0448 0.0268 Llama4Rec 0.0647* 0.0476* 0.0967* 0.0608* 0.0304* 0.0222* 0.0461* 0.0286* 0.0434* 0.0338* 0.057* 0.0394* Impro. 20.48% 24.93% 14.30% 19.92% 7.42% 9.36% 4.54% 8.33% 21.23% 24.26% 18.75% 22.36% MixGCF Base 0.0526 0.0401 0.0757 0.0496 0.0159 0.0115 0.0238 0.0147 0.0426 0.0330 0.0556 0.0384 IFT 0.0617 0.0452 0.0906 0.0570 0.0162 0.0114 0.0259 0.0154 0.0337 0.0243 0.0506 0.0312 Llama4Rec 0.0690* 0.0515* 0.0949* 0.0621* 0.0174* 0.0128* 0.0259 0.0162* 0.0495* 0.0384* 0.0635* 0.0441* Impro. 11.83% 13.94% 4.75% 8.95% 7.41% 11.30% 0.00% 5.19% 16.20% 16.36% 14.21% 14.84% SGL Base 0.0505 0.0380 0.0729 0.0472 0.0284 0.0206 0.0434 0.0267 0.0419 0.0319 0.0566 0.0380 IFT 0.0520 0.0392 0.0792 0.0503 0.0275 0.0202 0.0438 0.0269 0.0326 0.0237 0.0499 0.0307 Llama4Rec 0.0632* 0.0479* 0.0917* 0.0596* 0.0308* 0.0224* 0.0480* 0.0294* 0.0501* 0.0393* 0.0634* 0.0448* Impro. 21.54% 22.19% 15.78% 18.49% 8.45% 8.74% 9.59% 9.29% 19.57% 23.20% 12.01% 17.89%
Table 3. Performance achieved by different sequential recommendation methods.
Backbone Method ML-100K ML-1M BookCrossing H@3 $\uparrow$ N@3 $\uparrow$ H@5 $\uparrow$ N@5 $\uparrow$ H@3 $\uparrow$ N@3 $\uparrow$ H@5 $\uparrow$ N@5 $\uparrow$ H@3 $\uparrow$ N@3 $\uparrow$ H@5 $\uparrow$ N@5 $\uparrow$ SASRec Base 0.0187 0.0125 0.0385 0.0205 0.0277 0.0165 0.0502 0.0257 0.0086 0.0049 0.0163 0.0081 IFT 0.0204 0.0136 0.0379 0.0207 0.0241 0.0159 0.0473 0.0254 0.0124 0.0086 0.0185 0.0111 Llama4Rec 0.0238* 0.0155* 0.0449* 0.0240* 0.0293* 0.0201* 0.0504 0.0287* 0.0142* 0.0098* 0.0227* 0.0131* Impro. 16.67% 13.97% 16.62% 15.94% 5.78% 21.82% 0.40% 11.67% 14.52% 13.95% 22.70% 18.02% BERT4Rec Base 0.0153 0.0104 0.0294 0.0161 0.0107 0.0069 0.0211 0.0112 0.0088 0.0058 0.0161 0.0088 IFT 0.0174 0.0119 0.0326 0.0100 0.0106 0.0071 0.0188 0.0104 0.0127 0.0092 0.0180 0.0113 Llama4Rec 0.0198* 0.0134* 0.0332 0.0189* 0.0115* 0.0078* 0.0206 0.0115* 0.0154* 0.0108* 0.023* 0.0139* Impro. 13.79% 12.61% 1.84% 17.39% 7.48% 9.86% -2.37% 2.68% 21.26% 17.39% 27.78% 23.01% CL4SRec Base 0.0243 0.0143 0.0436 0.0222 0.0259 0.0153 0.0492 0.0248 0.0083 0.0048 0.0165 0.0082 IFT 0.0230 0.0149 0.0428 0.0230 0.0234 0.0155 0.0447 0.0241 0.0102 0.0071 0.0177 0.0102 Llama4Rec 0.0255* 0.0182* 0.0440 0.0255* 0.0278* 0.0185* 0.0482 0.0268* 0.0138* 0.0093* 0.0220* 0.0127* Impro. 4.94% 22.15% 0.92% 10.87% 7.34% 19.35% -2.03% 8.06% 35.29% 30.99% 24.29% 24.51%
5.1.3. Data Preprocessing.
Following the methodology of prior works (Zhang et al., 2023c; Luo et al., 2022b), we adopt a leave-one-out evaluation strategy. More specifically, within each user’s interaction sequence, we choose the most recent item as the test instance. The item immediately preceding this serves as the validation instance, while all remaining interactions are used to constitute the training set. Moreover, regarding the instruction-tuning dataset construction, we randomly sampled 5K instructions for each recommendation task on the ML-100K, ML-1M, and BookCrossing datasets, respectively. We eliminated instructions that were repetitive or of low quality (identified by users with fewer than three interactions in their interaction history), leaving approximately 25K high-quality instructions. These instructions are mixed to create an instruction-tuning dataset to fine-tune the LLM.
5.1.4. Backbone Models
We incorporate our Llama4Rec with the following recommendation models that are often used for various recommendation tasks as the backbone models:
  • Direct Recommendation. In the scenario of direct recommendation, We adopt four representative methods, including: MF (Koren et al., 2009), LightGCN (He et al., 2020), MixGCF (Huang et al., 2021), and SGL (Wu et al., 2021).
  • Sequential Recommendation. Regarding sequential recommendation, we opt for three widely used models, including: SASRec (Kang and McAuley, 2018), BERT4Rec (Sun et al., 2019), and CL4SRec (Xie et al., 2022).
  • Rating Prediction. We consider the following classical models for rating prediction, including: DeepFM (Guo et al., 2017), NFM (He et al., 2017), DCN (Wang et al., 2017), AFM (Xiao et al., 2017), xDeepFM (Lian et al., 2018), and AutoInt (Song et al., 2019).
We employ the LLaMA-2 7B version as the backbone LLM across all experiments, unless specifically mentioned otherwise. Our primary comparison is with the standard Instruction Fine-Tuning (IFT) method adopted in TALLRec (Bao et al., 2023) and InstructRec (Zhang et al., 2023c). For the rating prediction task, LLaMA-2 with IFT is used to directly predict the rating. For the top-k recommendation task, the tuned LLM is used to re-rank the list predicted by the backbone model, in accordance with (Hou et al., 2023), referred to as listwise ranking. Besides, we also adopt LLM for predicting the rating for each item and sort by the predicted scores, referred to as pointwise ranking.
5.1.5. Implementation Details.
During training for LLaMA 2 (7B) with full-parameter tuning, we use a uniform learning rate of $2\times 10^{-5}$ and a context length of 2048, and we set the batch size as 16. Additionally, we use a cosine scheduler for three epochs in total with a 50-step warm-up period. To efficiently train the computationally intensive models, we simultaneously employ DeepSpeed training with ZeRO-3 stage (Rajbhandari et al., 2020) and flash attention (Dao et al., 2022). We trained the 7B model on 16 NVIDIA A800 80GB GPUs. For the inference stage, we employed the vLLM framework (Kwon et al., 2023) with greedy decoding, setting the temperature to 0. Only one GPU was utilized during the inference phase. We only evaluate the instruction-tuned LLaMA model for rating prediction task since it is not applicable for directly making top-$k$ recommendations.
[…]
Table 4. Performance achieved by different methods in rating prediction task.
Backbone Method ML-100K ML-1M BookCrossing RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ LLaMA IFT 1.2792 0.8940 1.2302 0.8770 2.0152 1.3782 DeepFM Base 1.0487 0.8082 0.9455 0.7409 1.7738 1.3554 Llama4Rec 1.0306* 0.7987* 0.9360* 0.7321* 1.6958* 1.2843* Impro. 1.73% 1.18% 1.00% 1.19% 4.40% 5.25% NFM Base 1.0284 0.8005 0.9438 0.7364 2.121 1.5984 Llama4Rec 1.0189* 0.7961 0.9369* 0.7303* 1.9253* 1.4473* Impro. 0.92% 0.55% 0.73% 0.83% 9.23% 9.45% DCN Base 1.0478 0.8063 0.9426 0.7342 2.0216 1.4622 Llama4Rec 1.0367* 0.8033 0.9345* 0.7272* 1.8518* 1.3566* Impro. 1.06% 0.37% 0.86% 0.95% 8.40% 7.22% AFM Base 1.0471 0.8035 0.9508 0.7464 1.6516 1.2614 Llama4Rec 1.0340* 0.7996 0.9426* 0.7394* 1.6244* 1.2259* Impro. 1.25% 0.49% 0.86% 0.94% 1.65% 2.81% xDeepFM Base 1.1472 0.8836 0.9519 0.7428 2.1756 1.6461 Llama4Rec 1.0947* 0.8483* 0.9401* 0.7336* 1.9610* 1.4833* Impro. 4.58% 4.00% 1.24% 1.24% 9.86% 9.89% AutoInt Base 1.0500 0.8120 0.9471 0.7404 1.9148 1.4501 Llama4Rec 1.0369* 0.8059* 0.9382* 0.7326* 1.7917* 1.3492* Impro. 1.25% 0.75% 0.94% 1.05% 6.43% 6.96%
5.2. Main Results (RQ1)
We conducted an extensive evaluation of our proposed Llama4Rec and the baseline methods on three datasets to assess the model’s performance under diverse recommendation scenarios. The experiment results for rating prediction, direct recommendation, and sequential recommendation are shown in Table 2, Table 3, and Table 4, respectively. We have the following key observations.
  • Llama4Rec consistently outperforms baseline methods in almost all scenarios, with particularly significant improvements observed in the direct recommendation task. Moreover, our findings reveal that direct instruction fine-tuning LLMs for recommendation tasks does not consistently yield promising performance. These results highlight the effectiveness of integrating LLMs into conventional recommendation models, underscoring the importance of incorporating the mechanism that utilizes instruction-tuned LLM to mutually augment and adaptively aggregate with conventional recommendation models.
  • In the scenario of the rating prediction task, while the instruction-tuned LLaMA model significantly underperforms when compared to conventional recommendation models, integrating the LLM yields a marked performance improvement. This suggests that the LLM and conventional recommendation models learn distinct aspects of information. Consequently, integrating the LLM with conventional recommendation models could enhance recommendation performance.
  • In the context of top-$k$ recommendations, Llama4Rec exhibits a more pronounced improvement for direct recommendations task. In addition, a more significant enhancement is observed on the Bookcrossing dataset, which can be attributed to the more fine-grained and distinguishable rating of the Bookcrossing dataset.
Table 5. Ablation study on key components of Llama4Rec on the ML-1M dataset.
Models ML-1M BookCrossing H@3 $\uparrow$ N@3 $\uparrow$ H@3 $\uparrow$ N@3 $\uparrow$ LightGCN 0.0283 (-) 0.0203 (-) 0.0358 (-) 0.0272 (-) IFT 0.0268 (-5.30%) 0.0193 (-4.93%) 0.0287 (-19.84%) 0.0202 (-25.74%) Llama4Rec w/o DA 0.0294 (+3.89%) 0.0209 (+2.96%) 0.0408 (+13.97%) 0.0319 (+17.28%) Llama4Rec w/o PA 0.0277 (-2.12%) 0.0199 (-1.97%) 0.0372 (+3.92%) 0.0279 (+2.57%) Llama4Rec w/o AA 0.0298 (+5.30%) 0.0218 (+7.39%) 0.0429 (+19.83%) 0.0332 (+22.06%) Llama4Rec 0.0304 (+7.42%) 0.0222 (+9.36%) 0.0434 (+21.23%) 0.0338 (+24.26%)
5.3. Ablation Study (RQ2)
We conducted an ablation study to analyze the contributions of different components in our model. Table 5 summarizes the results of the ablation studies across three variants on the ML-1M dataset. It is evident that the full model performs considerably better than all its variants, indicating that all the main components contribute significantly to overall performance improvement. Moreover, compared to the conventional model, the instruction-tuned LLM does not achieve superior results, underscoring the importance of model aggregation. We further analyze the specific impact of each component, and our observations are as follows:
  • w/o Data Augmentation (w/o DA): In this variant, we remove the data augmentation module while maintaining other components the same. Experimental results reveal a obvious decline in performance when this module is excluded. This indicates the module’s capacity to mitigate data sparsity and long-tail problem, consequently enhancing model performance.
  • w/o Prompt Augmentation (w/o PA): In this variant, we remove the prompt augmentation component, a crucial element of the proposed framework. Experimental results demonstrate a significant degradation in model performance when this module is excluded, thereby validating its essential role. By employing the instruction-tuned LLM with prompt augmentation from prior knowledge by conventional recommendation models, we achieve an enhanced model performance, attributable to the capture of different aspects of information.
  • w/o Adaptive Aggregation (w/o AA): In this variant, we substitute adaptive aggregation with uniform aggregation and keep other modules unchanged. Experimental results demonstrate a drop in model performance, underscoring the significance of accounting for the user’s long-tail coefficient and employing adaptive aggregation.
5.4. Hyper-parameter Study (RQ3)
5.4.1. Analysis of Hyper-parameters $\alpha_{1}$ and $\alpha_{2}$
We conducted an analysis of the effects of hyper-parameters $\alpha_{1}$ and $\alpha_{2}$. These parameters play crucial roles in controlling the weight in adaptive aggregation, as illustrated in Equation (7). Figure 3 presents the results on the ML-1M dataset using LightGCN as the backbone model. As $\alpha_{1}$ increases, we observe an initial surge in the model’s performance, followed by a decline. This trend suggests appropriate selection of $\alpha_{1}$ would enhance the model performance. With respect to $\alpha_{2}$, we observe a similar trend but the decline is more pronounced. This observation is consistent with the principle of adaptive aggregation, which emphasizes the importance of assigning suitable weights to tail users.
Figure 3. Impact of hyper-parameters $\alpha_{1}$ and $\alpha_{2}$ on ML-1M dataset with backbone model LightGCN.
Figure 4. Performance comparison w.r.t different LLaMA-2 size for training Llama4Rec on the Bookcrossing dataset.
Figure 5. Performance comparison w.r.t different numbers of instructions for training Llama4Rec on the ML-1M dataset.
5.4.2. Analysis of Model Scaling.
We further instruction-tuned the LLaMA-2 model with different model size. [^5] Due to resource constraints, training the LLaMA-2 (70B) model with identical experimental settings was unfeasible, consistently leading to Out-Of-Memory (OOM) errors. A comparative analysis was conducted between the 7B and 13B variants of the instruction-tuned models, with performance differences specifically evaluated across various backbone models within the Bookcrossing dataset, as depicted in Figure 4. Our findings suggest that the LLaMA-2 (13B) model generally surpasses the 7B version in performance. This can be attributed to the superior language comprehension and reasoning abilities of the larger model, which contribute to improved recommendation results. However, it’s worth noting that the improvements are not substantial, indicating that while larger models may provide some performance benefits, the degree of improvement may not always justify the increased computational resources and training time required. It underscores the importance of considering the trade-off between model size, performance gain, and resource efficiency in the design and application of large language models.
5.4.3. Analysis of Data Scaling.
We evaluated the effect of data size on LLM training by varying the number of instructions in the instruction-tuning dataset. Proportionality with our original configuration, the model with 2.5K instructions underwent 250 training steps, while the 12.5K instructions version was trained over 1250 steps. As depicted in Figure 5, a clear trend emerges: model performance improves with an increase in the number of instructions, particularly for direct recommendation models. This highlights the importance of utilizing larger and more diverse datasets for instruction tuning LLMs to optimize performance.
5.5. Further Discussion
In this part, we discuss about the computational efficiency and future improvements. In the Llama4Rec framework, additional training with augmented data is required, which may present a potential limitation. In the current experimental setup, we train a new model from scratch. However, this process could be optimized by continuing to train a previously tuned model, thereby reducing time costs. Additionally, in our experiment, we observed that training the LLaMA-2 7B model with around 25K instructions on 16 A800 GPUs with 2500 steps took approximately 1.94 hours. The inference time for each instruction averaged about 17 instructions per second, translating to a requirement of around 0.059 seconds per item for computation by a single A800 GPU.
This training and inference duration significantly exceeds that of conventional recommendation models, highlighting the limitations of current LLM-based recommender systems. The substantial demand for computational resources also represents a significant challenge. Consequently, employing instruction LLMs for large-scale industrial recommender systems, such as those with millions of users, is presently impractical. However, future advancements in accelerated and parallel computing algorithms for language model inference could potentially reduce inference times and computation resources. This improvement might make the integration of LLMs into large-scale recommender systems feasible, especially by leveraging many GPUs for parallel computation.
6. Conclusion and future work
In this study, we present Llama4Rec, a general and model-agnostic framework tailored to facilitate mutual augmentation between conventional recommendation models and LLMs through data augmentation and prompt augmentation. Data augmentation for conventional recommendation models could alleviate issues of data sparsity and the long-tail problem, thus improving conventional recommendation model performance. Prompt augmentation, on the other hand, allows the LLM to externalize additional collaborative or sequential information and further enhance the model capability. Furthermore, adaptive aggregation is employed to merge the predictions from both kinds of augmented models, resulting in more optimized recommendation performance. Comprehensive experimental results across three diverse recommendation tasks on three real-world datasets demonstrate the effectiveness of Llama4Rec. While our current approach focuses on mutual augmentation within a single step, our future work will explore expanding mutual augmentation in an iterative manner, potentially unlocking further improvements in model performance.

### v2

@@
 Integrating Large Language Models into Recommendation via Mutual Augmentation and Adaptive Aggregation
-Abstract.
-Conventional recommendation methods have achieved notable advancements by harnessing collaborative or sequential information from user behavior.
-Recently, large language models (LLMs) have gained prominence for their capabilities in understanding and reasoning over textual semantics, and have found utility in various domains, including recommendation.
-Conventional recommendation methods and LLMs each have their own strengths and weaknesses.
-While conventional methods excel at mining collaborative information and modeling sequential behavior, they struggle with data sparsity and the long-tail problem.
-LLMs, on the other hand, are proficient at utilizing rich textual contexts but face challenges in mining collaborative or sequential information.
-Despite their individual successes, there is a significant gap in leveraging their combined potential to enhance recommendation performance.
-In this paper, we introduce a general and model-agnostic framework known as Large Language model with mutual augmentation and adaptive aggregation for Recommendation (Llama4Rec).
-Llama4Rec synergistically integrates conventional and LLM-based recommendation models.
-Llama4Rec proposes data augmentation and prompt augmentation strategies tailored to enhance the conventional model and the LLM respectively.
+Abstract
+Conventional recommender systems and Large Language Model (LLM)-based recommender systems each have their strengths and weaknesses.
+While conventional recommendation methods excel at mining collaborative information and modeling sequential behavior, they struggle with data sparsity and the long-tail problem.
+LLM, on the other hand, is proficient at utilizing rich textual contexts but faces challenges in mining collaborative or sequential information.
+Despite their individual successes, there is a significant gap in leveraging their ensemble potential to enhance recommendation performance.
+In this paper, we introduce a general and model-agnostic framework known as Large language models with mutual augmentation and adaptive aggregation for Recommendation (Llama4Rec), aiming to bridge this gap via explicitly ensemble LLM and conventional recommendation model for more effective recommendation.
+We propose data augmentation and prompt augmentation strategies tailored to enhance the conventional recommendation model and LLM respectively.
 An adaptive aggregation module is adopted to combine the predictions of both kinds of models to refine the final recommendation results.
-Empirical studies on three real-world datasets validate the superiority of Llama4Rec, demonstrating its consistent and significant improvements in recommendation performance over baseline methods.
-${}^{\dagger}$oresponding Author
-copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: The 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14-18, 2024; Washington D.C., USA
-1. Introduction
-Recommender systems have emerged as crucial solutions for mitigating the challenge of information overload (He et al., 2023a, b; Luo et al., 2023a, 2022a).
-Recommender systems encompass a multitude of tasks, such as rating prediction (Steck, 2013; Khan et al., 2021) and top-$k$ recommendation (Le and Lauw, 2021; Luo et al., 2023b).
-The top-$k$ recommendation, which encompasses collaborative filtering-based direct recommendation (He et al., 2017, 2020), sequential recommendation (Kang and McAuley, 2018; Sun et al., 2019; Chen et al., 2023), and more, has found wide applications in various areas.
+Empirical studies on three datasets validate the superiority of Llama4Rec, demonstrating significant improvements in recommendation performance.
+I Introduction
+Recommender systems have emerged as crucial solutions for mitigating the challenge of information overload [68, 37, 35, 38, 39, 32].
+Recommender systems encompass a multitude of tasks, such as rating prediction [48, 21, 36] and top-$k$ recommendation [24, 64, 34].
+The top-$k$ recommendation, which encompasses collaborative filtering-based direct recommendation [15, 14], sequential recommendation [20, 49, 4], and more, has found wide applications in various areas.
 However, recommender systems still suffer from the data sparsity and long-tail problem.
 Data sparsity arises from sparse user-item interactions, making the task of accurately capturing user preferences more challenging.
-The long-tail problem further intensifies data sparsity issue, as a substantial number of less popular items (i.e., long-tail items) are infrequently interacted with, leading to inadequate data for effective model training and compromised recommendation quality.
-In recent years, Large Language Models (LLMs) have emerged, exhibiting exceptional capabilities in language understanding, text generation, and complex reasoning tasks (OpenAI, 2023; Touvron et al., 2023a, b; Wang et al., 2023; Zhou et al., 2023).
-Recent studies have started exploring their applicability in recommender systems (Liu et al., 2023; Bao et al., 2023; Zhang et al., 2023c).
+The long-tail problem further intensifies the data sparsity issue, as a substantial number of less popular items (i.e., long-tail items) are infrequently interacted with, leading to inadequate data for effective model training and compromised recommendation quality.
+Figure 1: An example that illustrates the motivation of Llama4Rec.
+In recent years, Large Language Models (LLMs) have emerged, exhibiting exceptional capabilities in language understanding, text generation, and complex reasoning tasks [41, 51, 52, 2].
+Recent studies have started exploring their applicability in recommender systems [29, 3, 66].
 For example, Liu et al.
-employed ChatGPT with in-context learning (ICL) for various recommendation tasks (Liu et al., 2023).
-Further progress has been achieved by adopting the instruction tuning technique (Ouyang et al., 2022; Longpre et al., 2023) to align general-purpose LLMs with recommendation tasks for improved performance (Zhang et al., 2023c; Bao et al., 2023).
-For instance, TALLRec (Bao et al., 2023) reformulates the recommendation problem as a binary classification task and introduces an effective instruction fine-tuning framework for adapting the LLaMA model (Touvron et al., 2023b).
+employed ChatGPT with in-context learning (ICL) for various recommendation tasks [29].
+Further progress has been achieved by adopting the instruction tuning technique [42, 31] to align general-purpose LLMs with recommendation tasks for improved performance [66, 3].
+For instance, TALLRec [3] reformulates the recommendation problem as a binary classification task and introduces an effective instruction fine-tuning framework for adapting the LLaMA model [52].
 However, these LLM-based recommendation methods may not perform optimally as they do not harness the collaborative or sequential information captured by conventional recommendation models.
 Conventional recommendation models and LLM-based recommendation methods each have their respective strengths and weaknesses.
+An illustrative example is shown in Figure 1.
 Conventional methods excel in mining collaborative information and modeling sequential behaviors, while LLMs are proficient in leveraging rich textual contexts.
-As such, the integration of LLMs into recommender systems presents a significant opportunity to amalgamate the advantages of both methodologies while circumventing their respective shortcomings.
-There have been initial attempts to harness the strengths of both conventional and LLM-based recommenders (Geng et al., 2022; Zhang et al., 2023b; Zheng et al., 2023; Wei et al., 2023).
-Some efforts have sought to integrate collaborative/sequential information by enabling LLMs to comprehend user/item ID information (Geng et al., 2022; Zhang et al., 2023b; Zheng et al., 2023).
+Model ensemble has emerged as a powerful technique in the field of reliable machine learning [8, 75].
+One of the key advantages of model ensemble is its ability to improve model performance.
+Each model in the ensemble may capture different aspects of the underlying data patterns, leading to a more comprehensive and robust problem understanding.
+By aggregating the predictions of these models, we can obtain a final prediction that is often more accurate and reliable than that of any single model.
+Inspired by the concept of model ensemble, the integration of LLMs into recommender systems presents a significant opportunity to amalgamate the advantages of both methods while circumventing their respective shortcomings.
+There have been initial attempts to harness the strengths of both conventional and LLM-based recommender systems [11, 69, 72, 57].
+Some efforts have sought to integrate collaborative or sequential information by enabling LLMs to comprehend user/item ID information [11, 69, 72].
 For instance, a concurrent study by Zhang et al.
-encoded the semantic embedding into the prompt (Zhang et al., 2023b) and send it to LLM.
-On the other hand, some research works have aimed to augment conventional models using LLMs via data or knowledge augmentation (Wei et al., 2023; Xi et al., 2023).
-LLMRec (Wei et al., 2023) enhances recommender systems by deploying LLMs to augment the interaction graph, thereby addressing the challenges of data sparsity and low-quality side information.
+encodes the semantic embedding into the prompt and sends it to LLM [69].
+Besides, Zheng et al.
+[72] proposes adapting large language models for recommendation by integrating collaborative semantics, combining contextual understanding with collaborative filtering to enhance recommendation performance.
+On the other hand, some research works have aimed to augment conventional models using LLMs via data or knowledge augmentation [57, 61].
+For example, LLMRec [57] enhances recommender systems by deploying LLMs to augment the interaction graph, thereby addressing the challenges of data sparsity and low-quality side information.
 However, existing methods have several limitations.
 Firstly, current methods lack generalizability.
-The strategy of integrating ID information proves challenging to generalize across different domains and necessitates additional training.
-The current data augmentation method is not universally applicable, as it only addresses a limited number of recommendation scenarios.
-Secondly, current research primarily focuses on the integration at the data-level (e.g., data augmentation) or model-level (e.g., make LLM understand ID semantics), leaving the result-level integration largely unexplored.
-Lastly, there is an absence of a comprehensive framework that combines and integrates these methods into a single construct.
+Current strategy of integrating ID information in LLMs proves challenging to generalize across different domains and necessitates additional training.
+Moreover, the current data augmentation method is not universally applicable, as it only addresses a limited number of recommendation scenarios.
+Secondly, current research primarily focuses on the integration at the data-level (e.g., data augmentation) or model-level (e.g., making LLM understand ID semantics), leaving the result-level integration largely unexplored.
+Lastly, there is an absence of a comprehensive framework that combines and integrates these methods into a holistic solution.
 In light of these limitations, our objective is to explore the integration of conventional recommendation models and LLM-based recommendation methods in depth to address the above limitations and enhance recommendation performance.
-In this paper, we introduce a general framework known as Large language model with mutual augmentation and adaptive aggregation for Recommendation, referred to as Llama4Rec, for brevity.
-The core idea of Llama4Rec is to allow conventional recommendation models and LLM-based recommendation models to mutually augment each other, followed by an adaptive aggregation of the augmented models to yield more optimized results.
-Specifically, Llama4Rec performs data augmentation for conventional recommendation models by leveraging instruction-tuned LLM to alleviate the data sparsity and long-tail problem.
+In this paper, we introduce a generic framework named Large language model with mutual augmentation and adaptive aggregation for Recommendation, referred to as Llama4Rec, for brevity.
+The core idea of Llama4Rec is to allow conventional recommendation models and LLM-based recommendation models to mutually augment each other, followed by an adaptive aggregation of the augmented models to merge the textual and collaborative modality and yield more optimized results.
+Specifically, Llama4Rec performs data augmentation for conventional recommendation models by leveraging instruction-tuned LLM to alleviate the data sparsity and long-tail problem and thus enrich the collaborative modality.
 The data augmentation is tailored with different strategies depending on the recommendation scenarios.
-Furthermore, we use conventional recommendation models to perform prompt augmentation for LLMs.
+Furthermore, we use conventional recommendation models to perform prompt augmentation for LLMs, enhancing the textual modality.
 The prompt augmentation includes enriching collaborative information from similar users and providing prior knowledge from the conventional recommendation model within the prompt.
@@
 We conduct empirical studies on three real-world datasets, encompassing three different recommendation tasks, to validate the superiority of our proposed method.
-The results consistently demonstrate its superior performance over baseline methods, highlighting notable improvements in recommendation performance.
+The results consistently showcase its superior performance compared to baseline methods, achieving on average 14.65% improvement in direct recommendation, a 14.21% improvement in sequential recommendation, and a 3.10% improvement in rating prediction.
 In a nutshell, the contributions of this work are threefold.
-  • We introduce Llama4Rec, a general and model-agnostic framework to integrate LLM into conventional recommendation models.
+  • We introduce Llama4Rec, a general and model-agnostic framework to integrate LLM into conventional recommendation models, merging the textual modality with collaborative modality.
 Llama4Rec performs the data augmentation for conventional models to alleviate the data sparsity problem and improve model performance.
-The prompt augmentation is applied to LLM for leveraging the information captured by the conventional models.
-  • Llama4Rec employs an adaptive aggregation approach to combine the prediction from the conventional recommendation model and LLM for improved recommendation performance via leveraging and merging the information by both kinds of models.
+The prompt augmentation is applied to LLM for leveraging the information captured by the collaborative-modal recommendation models.
+  • Llama4Rec employs an adaptive aggregation approach to combine the prediction from the conventional recommendation model and LLM for improved recommendation performance via leveraging and merging the information captured by such kinds of models working on different modalities.
   • To validate the effectiveness of Llama4Rec, we conduct extensive experiments on three real-world datasets across three diverse recommendation tasks.
 The empirical results demonstrate that Llama4Rec outperforms existing baselines, exhibiting notable improvements across multiple performance metrics.
-2. Related Work
-2.1. Conventional Recommendation Methods
-Conventional recommendation methods serve as the cornerstone for the contemporary landscape of recommender systems (Zhang et al., 2019).
+II Related Work
+II-A Conventional Recommendation Methods
+Conventional recommendation methods serve as the cornerstone for the contemporary landscape of recommender systems [68].
 Representative recommendation tasks include rating prediction, collaborative filtering-based direct recommendation, and sequential recommendation, where the latter two are usually formulated as top-$k$ recommendation problems.
-Specifically, one of the seminal techniques is the use of matrix factorization for rating prediction, popularized by methods such as Singular Value Decomposition (SVD) (Koren et al., 2009).
-Collaborative filtering (CF) is another commonly used technique for the recommender systems (Adomavicius and Tuzhilin, 2005).
-Recent advancements have evolved CF techniques into more complex neural network architectures and graph-based models (He et al., 2017; Wang et al., 2019) to enhance the model performance.
-Sequential recommendation models incorporate temporal patterns into the recommendation pipeline.
-Techniques such as recurrent neural networks have been adapted for this purpose (Hidasi et al., 2015).
-Recent research focuses on applying attention mechanisms to further refine these models, leading to a noteworthy boost in performance (Sun et al., 2019; Kang and McAuley, 2018).
-Although conventional recommendation techniques are well-suited for capturing latent information associated with users and items, they often require a substantial amount of user-item interactions to provide accurate recommendations, which limits their effectiveness in data sparse and long-tail scenarios (Park and Tuzhilin, 2008).
-Figure 1.
-(i) The overall framework architecture of the proposed Llama4Rec consists of two main components: mutual augmentation and adaptive aggregation.
-The mutual augmentation includes data augmentation and prompt augmentation.
-(ii) Illustration of the data augmentation process encompasses three diverse recommendation scenarios.
-(iii) The pipeline of the adaptive aggregation module, which merges the predictions from the conventional recommendation model and the LLM.
-2.2. Large Language Model for Recommendation
-LLMs have brought a paradigm shift in numerous areas of machine learning, including recommendation methods (Fan et al., 2023).
-One of the most compelling advantages of LLM-based recommendation methods is their capacity for contextual understanding and in context learning (Dong et al., 2022).
-Inspired by this, reference (Liu et al., 2023) utilized ChatGPT across diverse recommendation tasks and found it to be effective in specific contexts, underpinned by robust experiments and human evaluations.
-Similarly, Wang et al.
-(Wang and Lim, 2023) introduced a novel zero-shot technique for next-item recommendations, further substantiating the utility of LLMs in this arena.
-However, it is noteworthy that these methods do not consistently demonstrate a marked improvement over conventional recommendation algorithms, which is largely attributable to the inherent misalignment between the general-purpose capabilities of LLMs and the specialized requirements of the recommendation task.
-To address this issue, recent studies further attempt to instruct tuning the specific LLM to align with human preference (Bao et al., 2023; Zhang et al., 2023c).
-Typically, these approaches involve creating an instruction tuning dataset in line with recommendation tasks, which is then used to tune the LLM for recommendation.
-Such methodologies have demonstrated improved performance in generating more aligned and accurate recommendations.
-Nevertheless, while LLMs excel at capturing intricate textual patterns, they may encounter challenges in comprehensively encoding user and item collaborative or sequential information.
-Though some concurrent studies (Zhang et al., 2023b; Zheng et al., 2023) aim to address this gap, they often lack in terms of generalizability and comprehensibility.
-In response to this challenge, we propose a novel framework designed to mitigate this issue.
-3. Preliminary
-We consider a recommender system with a set of users, denoted $\mathcal{U}=\{u_{1},u_{2},\ldots,u_{n}\}$, and a set of items, denoted =$\mathcal{I}=\{i_{1},i_{2},\ldots,i_{m}\}$.
-The rating prediction task aims to estimate the unknown values of $r_{ui}$ in the user-item interaction matrix $R\in\mathbb{R}^{n\times m}$, where each entry $r_{ui}$ is the rating assigned by user $u$ to item $i$.
-Different from rating prediction, top-$k$ recommendation focuses on identifying a subset of items $\mathcal{S}_{u}\subset\mathcal{I}$ for each user $u$.
-The subset is chosen to maximize a user-specific utility $U(u,\mathcal{S})$ with the constraint $|\mathcal{S}|=k$, which can be formally expressed as:
-(1) $\mathcal{S}_{u}={\arg\max}_{\mathcal{S}\subset\mathcal{I},|\mathcal{S}|=k}U(u,\mathcal{S}).$
-In the context of LLM-based recommendation methods, let ${LLM}$ represent the original LLM.
-These kinds of methods first utilize prompt to interpret the recommendation task into natural language.
-The LLM-based recommendation for user $u$ with in-context learning is denoted by ${r}_{u}={LLM}(\mathcal{P}_{u})$ where $\mathcal{P}_{u}$ is the recommendation prompt for user $u$.
-The recommendation prompt could either ask LLM to predict a rating towards a target item, or rank candidate items derived by retrieval models in top-k recommendation.
-To instruction fine-tune LLM, a dedicated dataset $\mathcal{D}_{ins}$ consisting of various instructions is utilized.
-The resulting instruction-tuned LLM is denoted as ${LLM}^{\prime}$.
-Therefore, the recommendation process in the fine-tuned model can be succinctly represented as $r_{u}={LLM}^{\prime}(\mathcal{P}_{u})$.
-4. Methodology
-4.1. Overview
+Specifically, one of the seminal techniques is the use of matrix factorization for rating prediction, popularized by methods such as Singular Value Decomposition (SVD) [22].
+Collaborative filtering (CF) is another commonly used technique for the recommender systems [1].
+Recent advancements have evolved CF techniques into more complex neural network architectures and graph-based models [15, 56] to enhance the model performance.
+On the other hand, sequential recommendation models incorporate temporal patterns into the recommendation pipeline.
+Techniques such as recurrent neural networks have been adapted for this purpose [16].
+Recent research focuses on applying attention mechanisms to further refine these models, leading to a noteworthy boost in performance [49, 20].
+Although conventional recommendation techniques are well-suited for capturing latent information associated with users and items, they often require a substantial amount of user-item interactions to provide accurate recommendations, which limits their effectiveness in data sparse and long-tail scenarios [43].
+Figure 2: (i) The overall framework architecture of the proposed Llama4Rec consists of two main components: mutual augmentation and adaptive aggregation.
+The red dashed lines denoted the data augmentation and prompt augmentation process, respectively.
+(ii) Illustration of the data augmentation process, encompassing three diverse recommendation scenarios.
+(iii) The pipeline of the adaptive aggregation module, which includes the aggregation of predictions from both types of models.
+The symbols $u$, $i$, and $a$ represents user, item, and attributes, respectively.
+II-B Large Language Model for Recommendation
+Large Language Models (LLMs) have ushered in a paradigm shift across various domains, including recommender systems [70, 10].
+Their ability to perform contextual understanding and reasoning offers significant advantages for recommendation tasks [27, 7].
+Existing approaches leveraging LLMs in recommender systems can be broadly categorized into two groups: LLMs as recommenders and LLMs as enhancers.
+The first category involves adapting LLMs to directly serve as recommenders, typically through prompting [5], in-context learning [54, 29], or fine-tuning [3, 66, 33].
+For instance, Liu et al.
+[29] explored the use of ChatGPT across diverse recommendation tasks, demonstrating its effectiveness in specific contexts through rigorous experiments and human evaluations.
+Similarly, Bao et al.
+[3] proposed TALLRec, which fine-tunes LLMs using instruction tuning datasets tailored to recommendation tasks, aligning LLMs with recommendation objectives.
+The second category employs LLMs as enhancers to augment traditional recommender systems by generating or extracting semantic features, such as item descriptions or user profiles [58, 60, 53, 71, 45].
+For example, Wei et al.
+[58] introduced LLMRec, which integrates LLMs with graph augmentation to enrich item representations for recommendation.
+By leveraging the semantic capabilities of LLMs, these approaches enhance the performance of conventional recommenders, offering a complementary strategy to improve recommendation quality.
+In contrast to prior work, which typically adopts either the LLM as recommender or LLM as enhancer approach, our method uniquely combines the strengths of both paradigms.
+By integrating LLMs as both direct recommenders and enhancers within a unified framework, our approach leverages contextual understanding and semantic feature extraction to achieve superior recommendation performance.
+TABLE I: Summary of notations used in this work.
+Notation Definition
+$u,i$ Users and items
+$\hat{y}_{ui}$ Predicted score of user $u$ for item $i$
+$e_{u},e_{i}$ Embeddings of user $u$ and item $i$
+$\mathcal{D}$ Training dataset
+$\mathcal{A}$ Set of attributes
+$\mathcal{P}$ Prompt for the LLM
+$\mathcal{L}_{\text{BPR}}$ Bayesian personalized ranking (BPR) loss
+$\ell_{u}$ Long-tail coefficient for user $u$
+$N(u)$ Number of interactions for user $u$
+$U$ Utility score
+$\Theta$ Original parameters of the LLM
+$\alpha_{1},\alpha_{2}$ Hyperparameters
+III Methodology
+III-A Overview
 Figure 2 depicts the architecture of Llama4Rec, which consists of three components: data augmentation, prompt augmentation, and adaptive aggregation.
-More specifically, Llama4Rec leverages an instruction-tuned LLM to enhance conventional recommendation systems through data augmentation.
-The specific data augmentation strategies for different recommendation situations are detailed in Section 4.2.
-In addition, we employ conventional recommendation models to augment the LLM via prompt augmentation, with details in Section 4.3.
-To further refine the predictions of the conventional model and the LLM, we propose a simple yet effective adaptive aggregation module in Section 4.4.
-Lastly, we describe the training strategy for LLM in Section 4.5.
-4.2. Data Augmentation for Conventional Recommendation Model
+More specifically, Llama4Rec leverages an instruction-tuned LLM to enhance conventional recommender systems through data augmentation.
+The specific data augmentation strategies for different recommendation situations are detailed in Section III-B.
+In addition, we employ conventional recommendation models to augment the LLM via prompt augmentation, with details in Section III-C.
+To further refine the predictions of the conventional model and the LLM, we propose a simple yet effective adaptive aggregation module in Section III-D.
+Lastly, we describe the training strategy for LLM in Section III-E.
+Summary of notations are listed in Table I.
+III-B Data Augmentation for Conventional Recommender System
 We design ad-hoc data augmentation strategies for different recommendation scenarios to mitigate prevalent issues of data sparsity and the long-tail problem.
@@
 In the context of direct recommendation, we capitalize on the power of the instruction-tuned LLM to predict items that a user may like or dislike.
-We form pairs of these items to calculate the Bayesian Personalized Ranking (BPR) (Rendle et al., 2012) loss.
+We form pairs of these items to calculate the Bayesian Personalized Ranking (BPR) [46] loss.
 For sequential recommendation, we harness the capabilities of the instruction-tuned LLM to predict items that are highly preferred by the user.
@@
 For rating prediction, we utilize the LLM to extract valuable side information (i.e., missing attributes), which is then seamlessly integrated as additional features within the training data.
-4.2.1. Data Augmentation for Direct Recommendation
-For direct recommendation, the Bayesian Personalized Ranking (BPR) loss is commonly used to optimize the model (Rendle et al., 2012).
+III-B1 Data Augmentation for Direct Recommendation
+For direct recommendation, the Bayesian Personalized Ranking (BPR) loss [46] is commonly used to optimize the model.
 The objective of BPR is to maximize the score difference between correctly recommended items and incorrectly recommended items, thereby improving the accuracy of recommendations.
 The BPR loss is defined as:
-(2) $\mathcal{L}_{BPR}=-\sum_{(u,i,j)\in\mathcal{D}}\log\sigma(\hat{y}_{ui}-\hat{y}_{uj}),$
+$\mathcal{L}_{BPR}=-\sum_{(u,i,j)\in\mathcal{D}}\log\sigma(\hat{y}_{ui}-\hat{y}_{uj}),$ (1)
 where $(u,i,j)$ refers to a triple of user-item pairs, and the user $u$ has interacted with item $i$ (positive item) and item $j$ (negative item).
@@
 $\hat{y}_{ui}$ denotes the predicted score or preference of user $u$ for item $i$.
-[…]
-(3) $\mathcal{L}_{BPR}^{\prime}=-\sum_{(u,i,j)\in\mathcal{D}^{\prime}}\log\sigma(\hat{y}_{ui}-\hat{y}_{uj}).$
-This data augmentation strategy leverages the power of the instruction tuned LLM to enhance the performance of the direct recommendation model.
-4.2.2. Data Augmentation for Sequential Recommendation
-[…]
+$\sigma$ is the sigmoid function and $\sigma(\hat{y}{ui}-\hat{y}{uj})$
+Inspired by this, we propose a data augmentation strategy that harnesses the capabilities of the LLM to enhance the direct recommendation.
+Initially, we randomly select pairs of items for a user $u$ and prompt the LLM to rank each pair based on the user’s likely preference.
+The ranking prediction based on LLM is then combined with the original data and used to train a direct recommendation model.
+Formally, let $(i_{u1},i_{u2})$ denote a pair of un-interacted items for a user $u$.
+The LLM is prompted to rank these items, denoted as $i^{+},i^{-}=\textit{LLM}(\mathcal{P}_{1})$, where $\mathcal{P}_{1}$ is the corresponding prompt and $i^{+}$ is the item preferred over $i^{-}$.
+The training data $\mathcal{D}$ is updated as $\mathcal{D}^{\prime}=\mathcal{D}\cup(u,i^{+},i^{-})$.
+The BPR loss is then updated as:
+$\mathcal{L}_{BPR}^{\prime}=-\sum_{(u,i,j)\in\mathcal{D}^{\prime}}\log\sigma(\hat{y}_{ui}-\hat{y}_{uj}).$ (2)
+III-B2 Data Augmentation for Sequential Recommendation
+For sequential recommendation, the data augmentation strategy involves enriching the sequence of interacted items with additional items predicted by the LLM.
+Let’s consider a user $u$ with a corresponding sequence of interacted items $\{i_{1},...,i_{l}\}$.
+We randomly sample a list of un-interacted items $\{i_{u1},...,i_{ut}\}$, and adopt the prompt $\mathcal{P}_{2}$ to ask the LLM to predict the item most likely to be preferred by the user, denoted as $i_{p}=\textit{LLM}(\mathcal{P}_{2})$.
+This predicted item $i_{p}$ is then randomly inserted into the user’s sequence, resulting in an augmented sequence $\{i_{1},...,i_{p},...,i_{l}\}$.
+This sequence is used to train a more powerful sequential recommendation model.
 By including additional items predicted by the LLM, we can enrich the sequence of items for each user, providing a more comprehensive representation of the user’s preferences.
 This, in turn, can enhance the performance of the conventional recommendation model, leading to more accurate recommendations.
-4.2.3. Data Augmentation for Rating Prediction
+III-B3 Data Augmentation for Rating Prediction
 In rating prediction tasks, we introduce the use of in-context learning (ICL) in LLMs to provide side information.
 This is primarily due to the fact that recommendation datasets may contain incomplete information.
-For instance, the popular Movielens dataset (Harper and Konstan, 2015) lacks information about the director of the movies, which can hinder the performance of a conventional rating prediction model.
+For instance, the popular Movielens dataset [13] lacks information about the director of the movies, which can hinder the performance of a conventional rating prediction model.
 To mitigate this issue, we leverage the extensive world knowledge contained in an LLM.
 We prompt the LLM to provide side information, acting as additional attributes for users/items.
-[…]
-Top-k Recommendation Prompt Example:
-Instruction: Rank the candidate movies based on user historical interactions and make the top k recommendations.
-Interaction History: Beyond Rangoon (1995); Alien (1979); Hollow Reed (1996); Primary Colors (1998); …; Birds, The (1963)
-Candidate Items: Last Dance (1996); Remains of the Day, The (1993); Assassins (1995); …; Fatal Instinct (1993)
-Similar User Interaction History: L.A. Confidential (1997); Apt Pupil (1998); Kolya (1996); …; Star Wars (1977)
-Conventional Model Prediction: Remains of the Day, The (1993); Addiction, The (1995); …; Fugitive, The (1993)
-Output: Fugitive, The (1993); Angel Baby (1995); …; Remains of the Day, The (1993)
-Rating Prediction Prompt Example:
-Instruction: Predict the rating of a target movie based on the user’s historical movie ratings.
-Rating History: Independence Day (1996): 3; Grosse Fatigue (1994): 3; Face/Off (1997): 4; …; Shall We Dance? (1996): 3
-Candidate Item: Pink Floyd - The Wall (1982)
-Similar User Rating History: L.A. Confidential (1997): 3; Apt Pupil (1998): 4; …; English Patient, The (1996): 3
-Conventional Model Prediction: 3.2
-Output: 3
-Figure 2.
-Examples of instructions for top-k recommendation and rating prediction.
-The prompt augmentation component is underlined.
-To improve readability and facilitate better experimental evaluation, we introduce certain modifications to the original instructions employed in our experiments.
-4.3. Prompt Augmentation for Large Language Model
-Previous works (Bao et al., 2023; Zhang et al., 2023c) instruction tuning LLM for recommendation in a standard manner.
+Formally, we denote the rating prediction model as $\mathcal{M}_{r}$, and the attribute set as $\mathcal{A}=\{a_{1},a_{2},...,a_{n}\}$, where $a_{i}\in\mathcal{A}$ denotes a distinct attribute.
+The model predicts the rating as $\mathtt{Pred}=\mathcal{M}_{r}(\mathcal{A})$.
+We then prompt the LLM to provide additional attributes, where the prompt $\mathcal{P}_{3}$ contains some corresponding examples followed by detailed instructions.
+The process is denoted as $\{a_{d1},a_{d2},...\}=\textit{LLM}(\mathcal{P}_{3})$.
+The augmented attribute set is then formed as $\mathcal{A}^{\prime}=\mathcal{A}\cup\{a_{d1},a_{d2},$...\}.
+The model then predicts the rating using the augmented attribute set, denoted as $\mathtt{Pred}^{\prime}=\mathcal{M}_{r}(\mathcal{A}^{\prime})$.
+This approach allows us to leverage the LLM’s world knowledge to enhance the performance of the rating prediction model.
+III-C Prompt Augmentation for LLM
+Previous works [3, 66] instruction tuning LLM for recommendation in a standard manner.
 However, these methods can be sub-optimal due to the challenges of distinguishing users based solely on text-based prompt descriptions.
-Although some concurrent studies (Zhang et al., 2023b; Zheng et al., 2023) incorporate unique identifiers to differentiate users, these approaches require complex semantic understanding of IDs and additional training, limiting their generalizability.
-In this section, we introduce two text-based prompt augmentation strategies for LLM-based recommendations, i.e., we incorporate additional information within the prompt to enhance the model performance.
+Although some concurrent studies [69, 72] incorporate unique identifiers to differentiate users, these approaches require complex semantic understanding of IDs and additional training, limiting their generalizability.
+We introduce two text-based prompt augmentation strategies for LLM-based recommendations, i.e., we incorporate additional information within the prompt to enhance the model performance.
 First, we propose prompt augmentation with similar users, identifying users with analogous preferences to enrich the prompt, thereby enhancing the LLM’s ability to leverage collaborative information and generate personalized recommendations.
-Second, we propose prompt augmentation with conventional model prediction, providing prior knowledge to guide the LLM toward recommendations that align with user preferences.
+Second, we propose prompt augmentation with conventional recommendation model prediction, providing prior knowledge to guide the LLM toward recommendations that align with user preferences.
 Collectively, these strategies harness the strengths of both LLMs and conventional recommendation models, ensuring generalizability across a wide range of recommendation scenarios.
 The illustration of prompt augmentation is underlined in Figure 3.
-4.3.1. Prompt Augmentation with Collaborative Information from Similar User
-To incorporate collaborative information within the prompt and facilitate LLM reasoning, we introduce a prompt augmentation strategy with similar user.
+III-C1 Prompt Augmentation with Collaborative Information from Similar Users
+To incorporate collaborative information within the prompt and facilitate LLM reasoning, we introduce a prompt augmentation strategy with similar users.
 Initially, we utilize a pre-trained conventional recommendation model to acquire embeddings for each user.
-These embeddings represent users in a latent space, which encapsulates their preferences and behaviors.
+These embeddings serve as a representation of users in a latent space, which encapsulates their preferences and behaviors.
 Specifically, for a user $u$, in conjunction with a conventional recommendation model $\mathcal{M}_{c}$, we use $\mathcal{M}_{c}$ to obtain embeddings for each user, denoted as $\{e_{1},...,e_{n}\}$.
-These embeddings encapsulate the preferences and behaviors of the users, serving as a compact representation of the users in a latent space.
 We then calculate the similarity between these embeddings in the latent space.
@@
 In this paper, we calculate the cosine similarity to measure how closely two vectors align, denoted as:
-(4) $sim(u,v)=\frac{e_{u}\cdot e_{v}}{||e_{u}||\cdot||e_{v}||},$
+[…]
 where $e_{u}$ and $e_{v}$ are the embeddings of user $u$ and $v$, respectively, and $u,v\in\mathcal{U}$.
-The |||||$\cdot$||| | ⋅ | | denotes the Euclidean norm and $\cdot$ denotes the dot product.
+The $||\cdot||$ denotes the Euclidean norm and $\cdot$ denotes the dot product.
 We identify the pair of users $(u,v)$ that have the highest similarity, indicating that they are the most similar in terms of their preferences and behaviors.
@@
 This strategy leverages the collaborative information gleaned from similar users to generate more relevant and accurate prompts, thereby enhancing the recommendation performance of the LLM.
-4.3.2. Prompt Augmentation with Prior Knowledge from Conventional Recommendation Model Prediction
+III-C2 Prompt Augmentation with Prior Knowledge from Conventional Recommendation Model Prediction
 To enable the LLM to leverage information captured by conventional models, we propose a prompt augmentation method that incorporates information from conventional recommendation models.
@@
 Through augmenting prompts with predictions from conventional recommendation models, our method integrates collaborative or sequential information captured by these models, thereby enhancing the LLM’s contextual understanding and reasoning capabilities and resulting in better recommendation performance.
-Notably, unlike ID-based methods such as (Zhang et al., 2023b; Zheng et al., 2023), our approach relies entirely on text, enabling easy adaptation to new situations.
-Also, the prompt augmentation could be used as a plug-and-play component for recommendation with closed source LLM, such as the GPT-4 model (OpenAI, 2023).
-4.4. Adaptive Aggregation
-We endeavor to aggregate the outputs of LLM and conventional recommendation models at the result level for improved performance, considering the disparate model structures.
-However, indiscriminate aggregation of model predictions can potentially lead to suboptimal results.
+Notably, unlike ID-based methods such as [69, 72], our approach relies entirely on text, enabling easy adaptation to new situations.
+Moreover, the prompt augmentation could be used as a plug-and-play component for recommendation with closed source LLM, such as the GPT-4 model [41].
+III-D Adaptive Aggregation
+We endeavor to merge the textual and collaborative modality of both the LLM and conventional recommendation models at the post-learning stage.
+Considering the disparate model structures, we aggregate the outputs from both types of models for improving recommendation performance.
+However, indiscriminate aggregation of model predictions can lead to sub-optimal results.
 Conventional recommendation models, known for their susceptibility to the long tail issue, often struggle when dealing with the tail segment.
@@
 We first define the long-tail coefficient $\ell_{u}$ for user $u$ to quantify where the user is located in the tail of the distribution. The long-tail coefficient is defined as follows:
-(5) $\ell_{u}=\text{log}\left(N({u})+1\right),$
-where $N({u})$ is the number of interaction for user $u$. A lower long-tail coefficient value indicates that the user has fewer feedback.
+$\ell_{u}=\text{log}\left(N({u})+1\right),$ (4)
+where N(u)N({u}) is the number of interaction for user $u$. A lower long-tail coefficient value indicates that the user has fewer feedback.
 While the overarching architecture remains consistent, the implementation details are different for the two tasks considered, namely rating prediction and top-$k$ recommendation.
-4.4.1. Adaptive Aggregation for Rating Prediction
+III-D1 Adaptive Aggregation for Rating Prediction
 For the rating prediction task, we employ an instruction-tuned LLM to predict user-item utility scores directly.
-This approach incorporates the understanding of complex semantics and context by the LLM, which might be overlooked by traditional models.
+This approach incorporates the understanding of complex semantics and context by the LLM, which might be overlooked by conventional models.
 Similarly, conventional recommendation methods leverage collaborative information and user/item features for predicting the user rating.
-Specifically, the utility weight for user $u$, denoted as $U_{u}$, is directly set as its user rating.
-Subsequently, the LLM is engaged to predict the rating, symbolized as $U_{LLM}$.
-There are various methods to derive a final result based on the utility scores, such as training a neural network to process the utility scores from LLM and conventional models, yielding a final output via learning the complex reflection.
+Specifically, the LLM is adopted to predict the user rating as utility score, denoted as $U_{LLM}$.
+Besides, the conventional predicts the user rating as $U_{Rec}$.
+Subsequently, there are various methods to derive a final utility score based on both kinds of models, such as training a neural network to process the utility scores from LLM and conventional models, yielding a final output via learning a complex reflection.
 However, for the sake of simplicity in this paper, we adopt a simple yet effective linear interpolation approach.
-The final utility score for a user $u$ amalgamates the values from both models, represented as:
-(6) $U_{u}=\alpha_{u}{U}_{LLM}+(1-\alpha_{u})U_{Rec},$
+Thus, the final utility score for a user $u$ amalgamates the values from both models, represented as:
+$U_{u}=\alpha_{u}{U}_{LLM}+(1-\alpha_{u})U_{Rec},$ (5)
 where $\alpha_{u}$ is the adaptive parameter to control the weight for each model’s utility value for user $u$. We define the $\alpha_{u}$ as:
-(7) $\alpha_{u}=\text{max}\left(\frac{\ell_{max}-\ell_{u}}{\ell_{max}-\ell_{min}},\alpha_{2}\right)\cdot\alpha_{1},$
+$\alpha_{u}=\text{max}\left(\frac{\ell_{max}-\ell_{u}}{\ell_{max}-\ell_{min}},\alpha_{2}\right)\cdot\alpha_{1},$ (6)
 where $\ell_{max}$ and $\ell_{min}$ are the maximum and minimum long-tail coefficients of the users, respectively, $\alpha_{1}$ is a hyper-parameter that controls the weight, and $\alpha_{2}<1$ is a cut-off weight.
@@
 This aligns with the motivation we previously discussed.
-4.4.2. Adaptive Aggregation for Top-$k$ Recommendation
+III-D2 Adaptive Aggregation for Top-$k$ Recommendation
 For the top-$k$ recommendation task, the LLM is employed to re-rank the item list generated by a conventional recommendation model.
@@
 Then we assign utility scores for items at each position, denoted as $U_{1},{U}_{2},...,U_{k^{\prime}}$, where $U_{1}\geq U_{2}\geq...\geq U_{k^{\prime}}$.
-The final utility score for an item amalgamates the values from both the original rating and the LLM prediction, similar to the Equation (6).
-4.5. Training Strategy for LLM
-4.5.1. Instruction Tuning Dataset Construction
+In this paper, we also set ${U}_{p}=-p\cdot\mathcal{C}$ where ${p}\in\{1,2,...,{k^{\prime}}\}$, and Up{U}{p} is the utility score assigned to the item at position $p$.
+The final utility score for an item amalgamates the values from both the original rating and the LLM prediction, similar to Equation (5).
+III-E Training Strategy for LLM
+III-E1 Instruction Tuning Dataset Construction
 This section details the creation of an instruction-tuning dataset that encompasses two types of recommendation tasks catering to top-$k$ recommendation and rating prediction scenarios.
-A depiction of these two tasks, specifically referred to as listwise ranking and rating prediction, can be found in Figure 2.
+A depiction of these two tasks, specifically referred to as listwise ranking and rating prediction, can be found in the Figure 3.
 It is noteworthy that we also employ the LLM to execute pointwise ranking within top-$k$ recommendation scenarios, i.e., utilizing LLM to predict ratings for each item within the top-k recommendations and sorting the predicted ratings to derive the final result.
-4.5.2. Optimization via Instruction Tuning
+In our dataset, the input $x$ consists of user and item features (e.g., user preferences, item attributes) used to prompt the model, while the output $y$ represents the corresponding recommendation or prediction.
+The ground-truth information from the training set is used to construct the output $y$, ensuring that the model learns from accurate and verified data.
+Figure 3: Examples of instructions for top-k recommendation and rating prediction.
+The prompt augmentation component is highlighted.
+The similar user histories and conventional model predictions are prompt augmented, which are highlighted in light purple and beige.
+We keep the prompt template structure and convert the bullet-point format into cohesive sentences for use in experiments.
+In the top-k recommendation, both direct and sequential recommendations follow the same prompt structure, with the key difference being that the user interaction history for sequential recommendations is time-series aware.
+III-E2 Optimization via Instruction Tuning
 In this work, we perform full parameter instruction tuning to optimize LLMs using generated instruction data.
-Due to our need for customization, we chose LLaMA-2 (Touvron et al., 2023b), an open-source, high-performing LLM, which permits task-specific fine-tuning.
-During supervised fine-tuning, we apply a standard cross-entropy loss following Alpaca (Taori et al., 2023).
+Due to our need for customization, we chose LLaMA-2 [52], an open-source, high-performing LLM, which permits task-specific fine-tuning.
+During supervised fine-tuning, we apply a standard cross-entropy loss following Alpaca [50].
 The training set $\mathcal{D}_{ins}$ consists of instruction input-output pairs $(x,y)$, which have been represented in natural language.
 The objective is to fine-tune the pre-trained LLM by minimizing the cross-entropy loss, formalized as:
-(8) $\min_{\Theta}\sum_{(x,y)\in\mathcal{D}_{ins}}\sum_{t=1}^{|y|}-\log P_{\Theta}\left(y_{t}\mid x,y_{[1:t-1]}\right),$
+$\min_{\Theta}\sum_{(x,y)\in\mathcal{D}_{ins}}\sum_{t=1}^{|y|}-\log P_{\Theta}\left(y_{t}\mid x,y_{[1:t-1]}\right),$ (7)
 where $\Theta$ are the original parameters for LLM, $P_{\Theta}$ is the conditional probability, $|y|$ is the number of tokens in $y$, $y_{t}$ is the $t$-th token in the target output $y$, and $y_{[1:t-1]}$ represents tokens preceding $y_{t}$ in $y$.
-By minimizing this loss function, the model fine-tunes its parameters $\Theta$ to adapt to the specifics of the new instruction tuning dataset $\mathcal{D}_{ins}$, while leveraging the general language understanding and reasoning that has been acquired during pre-training (Zhang et al., 2023a).
+By minimizing this loss function, the model fine-tunes its parameters $\Theta$ to adapt to the specifics of the new instruction tuning dataset $\mathcal{D}_{ins}$, while leveraging the general language understanding and reasoning that has been acquired during pre-training [67].
 In this manner, LLM can capture the user’s preferences for items expressed in natural language, facilitating diverse recommendation tasks, including top-k recommendation and rating prediction.
-5. Experiment
-In this section, we present a thorough empirical evaluation to validate the effectiveness of our proposed framework.
-Specifically, our objective is to investigate whether the incorporation of our proposed Llama4Rec could enhance existing recommendation models.
-The overarching goal is to answer the following research questions:
-  • RQ1: Does our proposed Llama4Rec framework enhance the performance of existing recommendation models?
-  • RQ2: How do the various modules in Llama4Rec affect the recommendation performance?
-  • RQ3: How do different hyper-parameters impact the overall performance of the framework?
-5.1. Experiment Setup
-5.1.1. Dataset
-Following (Bao et al., 2023), we rigorously evaluate the performance of our proposed framework by employing three heterogeneous, real-world datasets.
-MovieLens [^1] https://grouplens.org/datasets/movielens/ (Harper and Konstan, 2015) serve as benchmark datasets in the realm of movie recommendation methods.
-We employ two variants of the dataset: MovieLens-100K (ML-100K) and MovieLens-1M (ML-1M).
-The former consists of approximately 100,000 user-item ratings, while the latter scales up to roughly 1,000,000 ratings.
-BookCrossing [^2] Due to the absence of timestamp data, we synthesize historical interactions through random sampling.
-(Ziegler et al., 2005) includes user-generated book ratings on a scale of 1 to 10, alongside metadata such as ‘Book-Author’ and ‘Book-Title’.
-We employ LLM to augment the ‘director’ and ‘star’ features for ML-100K and ML-1M datasets, and augment the ‘genre’ and ‘page length’ features for the BookCrossing dataset.
-To ensure the data quality, we adopt the 5-core setting, i.e., we filter unpopular users and items with fewer than five interactions for the BookCrossing dataset.
-The key characteristics of these datasets are delineated in Table 1.
-Table 1. Dataset Description.
-ML-100K ML-1M BookCrossing # of User 943 6,040 6,851 # of Item 1,682 3,706 9,085 # of Rating 100,000 1,000,209 115,219 Density 0.063046 0.044683 0.001851 User Features Gender, ZipCode, Gender, ZipCode, Location, Age Occupation, Age Occupation, Age Item Features Title, Genres Title, Genres Title, Author, Year Year, Publisher Augmented Movie Director, Movie Director, Book Genres, Features Movie Star Movie Star Page Length
-5.1.2. Evaluation Metrics.
-Aligning with (He et al., 2020; Sun et al., 2019), for the top-$k$ recommendation task, we turn to two well-established metrics: Hit Ratio (HR) and Normalized Discounted Cumulative Gain (NDCG), denoted by H and N, respectively.
-In our experiments, $k$ is configured to be either 3 or 5 for a comprehensive evaluation, similar to the experiment setting in (Zhang et al., 2023c).
-In accordance with (Fan et al., 2019), we employ Root Mean Squared Error (RMSE) and Mean Absolute Error (MAE) as evaluation metrics to ascertain the performance of the rating prediction task.
-Table 2. Performance achieved by different direct recommendation methods.
+TABLE II: Performance achieved by different direct recommendation methods.
+H@$k$ indicates the hit rate at position $k$, and N@$k$ represents normalized discounted cumulative gain (NDCG) at position $k$.
 Backbone Method ML-100K ML-1M BookCrossing H@3 $\uparrow$ N@3 $\uparrow$ H@5 $\uparrow$ N@5 $\uparrow$ H@3 $\uparrow$ N@3 $\uparrow$ H@5 $\uparrow$ N@5 $\uparrow$ H@3 $\uparrow$ N@3 $\uparrow$ H@5 $\uparrow$ N@5 $\uparrow$ MF Base 0.0455 0.0325 0.0690 0.0420 0.0255 0.0187 0.0403 0.0248 0.0294 0.0227 0.0394 0.0269 IFT 0.0546 0.0388 0.0790 0.0488 0.0242 0.0175 0.0410 0.0244 0.0247 0.0177 0.0377 0.0230 Llama4Rec 0.0645* 0.0474* 0.0919* 0.0588* 0.0281* 0.0203* 0.0433* 0.0265* 0.0365* 0.0284* 0.0462* 0.0324* Impro.
@@
 21.54% 22.19% 15.78% 18.49% 8.45% 8.74% 9.59% 9.29% 19.57% 23.20% 12.01% 17.89%
-Table 3. Performance achieved by different sequential recommendation methods.
+IV Experiment
+In this section, we present a thorough empirical evaluation to validate the effectiveness of our proposed framework.
+Specifically, our objective is to investigate whether the incorporation of our proposed Llama4Rec can enhance existing recommendation models.
+The overarching goal is to answer the following research questions:
+  • RQ1: Does our proposed Llama4Rec framework enhance the performance of existing recommendation models?
+  • RQ2: How do the various modules in Llama4Rec affect the recommendation performance?
+  • RQ3: How do different hyper-parameters impact the overall performance of the framework?
+RQ4: How does the instruction-tuned model compare to other LLMs, such as GPT?
+How would the model perform w.r.t.
+model and data scaling?
+IV-A Experiment Setup
+IV-A1 Dataset
+Following [3], we rigorously evaluate the performance of our proposed framework by employing three heterogeneous, real-world datasets.
+MovieLens [^1] https://grouplens.org/datasets/movielens/ [13] serve as benchmark datasets in the realm of movie recommendation.
+We employ two variants of the dataset: MovieLens-100K (ML-100K) and MovieLens-1M (ML-1M).
+The former consists of approximately 100,000 user-item ratings, while the latter scales up to roughly 1,000,000 ratings.
+BookCrossing [^2] Due to the absence of timestamp data, we synthesize historical interactions through random sampling.
+[74] includes user-generated book ratings on a scale of 1 to 10, alongside metadata such as ‘Book-Author’ and ‘Book-Title’.
+We employ the LLM to augment the ‘director’ and ‘star’ features for ML-100K and ML-1M datasets, and augment the ‘genre’ and ‘page length’ features for the BookCrossing dataset.
+To ensure the data quality, we adopt the 5-core setting, where we filter unpopular users and items with fewer than five interactions for the BookCrossing dataset.
+The key characteristics of these datasets are delineated in the Table III.
+TABLE III: Dataset Description.
+ML-100K ML-1M BookCrossing # of User 943 6,040 6,851 # of Item 1,682 3,706 9,085 # of Rating 100,000 1,000,209 115,219 Density 0.063046 0.044683 0.001851 User Features Gender, ZipCode, Gender, ZipCode, Location, Age Occupation, Age Occupation, Age Item Features Title, Genres Title, Genres Title, Author, Year Year, Publisher Augmented Movie Director, Movie Director, Book Genres, Features Movie Star Movie Star Page Length
+IV-A2 Evaluation Metrics.
+Aligning with [14, 49], for the top-$k$ recommendation task, we turn to two well-established metrics: Hit Ratio (HR) and Normalized Discounted Cumulative Gain (NDCG), denoted by H and N, respectively.
+In our experiments, $k$ is configured to be either 3 or 5 for a comprehensive evaluation, similar to the experiment setting in [66].
+For instance, H@$k$ refers to the hit rate at position $k$, while N@$k$ denotes the normalized discounted cumulative gain (NDCG) at position $k$.
+In accordance with [9], we employ Root Mean Squared Error (RMSE) and Mean Absolute Error (MAE) as evaluation metrics for the rating prediction task.
+TABLE IV: Performance achieved by different sequential recommendation methods.
+H@$k$ indicates the hit rate at position $k$, and N@$k$ represents normalized discounted cumulative gain (NDCG) at position $k$.
 Backbone Method ML-100K ML-1M BookCrossing H@3 $\uparrow$ N@3 $\uparrow$ H@5 $\uparrow$ N@5 $\uparrow$ H@3 $\uparrow$ N@3 $\uparrow$ H@5 $\uparrow$ N@5 $\uparrow$ H@3 $\uparrow$ N@3 $\uparrow$ H@5 $\uparrow$ N@5 $\uparrow$ SASRec Base 0.0187 0.0125 0.0385 0.0205 0.0277 0.0165 0.0502 0.0257 0.0086 0.0049 0.0163 0.0081 IFT 0.0204 0.0136 0.0379 0.0207 0.0241 0.0159 0.0473 0.0254 0.0124 0.0086 0.0185 0.0111 Llama4Rec 0.0238* 0.0155* 0.0449* 0.0240* 0.0293* 0.0201* 0.0504 0.0287* 0.0142* 0.0098* 0.0227* 0.0131* Impro.
@@
 4.94% 22.15% 0.92% 10.87% 7.34% 19.35% -2.03% 8.06% 35.29% 30.99% 24.29% 24.51%
-5.1.3. Data Preprocessing.
-Following the methodology of prior works (Zhang et al., 2023c; Luo et al., 2022b), we adopt a leave-one-out evaluation strategy.
+IV-A3 Data Preprocessing.
+Following prior works [66, 40], we adopt a leave-one-out evaluation strategy.
 More specifically, within each user’s interaction sequence, we choose the most recent item as the test instance.
 The item immediately preceding this serves as the validation instance, while all remaining interactions are used to constitute the training set.
-Moreover, regarding the instruction-tuning dataset construction, we randomly sampled 5K instructions for each recommendation task on the ML-100K, ML-1M, and BookCrossing datasets, respectively.
+Moreover, regarding the instruction-tuning dataset construction, we randomly sampled 5K instructions for each of the three recommendation tasks (rating prediction, direct recommendation, and sequential recommendation) across three datasets (ML-100K, ML-1M, and BookCrossing), resulting in a total of 3 × 3 × 5K = 45K raw data.
 We eliminated instructions that were repetitive or of low quality (identified by users with fewer than three interactions in their interaction history), leaving approximately 25K high-quality instructions.
 These instructions are mixed to create an instruction-tuning dataset to fine-tune the LLM.
-5.1.4. Backbone Models
+IV-A4 Backbone Models
 We incorporate our Llama4Rec with the following recommendation models that are often used for various recommendation tasks as the backbone models:
-  • Direct Recommendation.
-In the scenario of direct recommendation, We adopt four representative methods, including: MF (Koren et al., 2009), LightGCN (He et al., 2020), MixGCF (Huang et al., 2021), and SGL (Wu et al., 2021).
-  • Sequential Recommendation.
-Regarding sequential recommendation, we opt for three widely used models, including: SASRec (Kang and McAuley, 2018), BERT4Rec (Sun et al., 2019), and CL4SRec (Xie et al., 2022).
-  • Rating Prediction.
-We consider the following classical models for rating prediction, including: DeepFM (Guo et al., 2017), NFM (He et al., 2017), DCN (Wang et al., 2017), AFM (Xiao et al., 2017), xDeepFM (Lian et al., 2018), and AutoInt (Song et al., 2019).
+  • Direct Recommendation. In the scenario of direct recommendation, We adopt four representative methods, including: MF [22], LightGCN [14], MixGCF [19], and SGL [59].
+  • Sequential Recommendation. Regarding sequential recommendation, we opt for three widely used models, including: SASRec [20], BERT4Rec [49], and CL4SRec [63].
+  • Rating Prediction. We consider the following classical models for rating prediction, including: DeepFM [12], NFM [15], DCN [55], AFM [62], xDeepFM [26], and AutoInt [47].
 We employ the LLaMA-2 7B version as the backbone LLM across all experiments, unless specifically mentioned otherwise.
-Our primary comparison is with the standard Instruction Fine-Tuning (IFT) method adopted in TALLRec (Bao et al., 2023) and InstructRec (Zhang et al., 2023c).
+Our primary comparison is with the standard Instruction Fine-Tuning (IFT) method adopted in TALLRec [3] and InstructRec [66].
 For the rating prediction task, LLaMA-2 with IFT is used to directly predict the rating.
-For the top-k recommendation task, the tuned LLM is used to re-rank the list predicted by the backbone model, in accordance with (Hou et al., 2023), referred to as listwise ranking.
+For the top-k recommendation task, the tuned LLM is used to re-rank the list predicted by the backbone model, in accordance with [18], referred to as listwise ranking.
 Besides, we also adopt LLM for predicting the rating for each item and sort by the predicted scores, referred to as pointwise ranking.
-5.1.5. Implementation Details.
+IV-A5 Implementation Details.
 During training for LLaMA 2 (7B) with full-parameter tuning, we use a uniform learning rate of $2\times 10^{-5}$ and a context length of 2048, and we set the batch size as 16.
 Additionally, we use a cosine scheduler for three epochs in total with a 50-step warm-up period.
-To efficiently train the computationally intensive models, we simultaneously employ DeepSpeed training with ZeRO-3 stage (Rajbhandari et al., 2020) and flash attention (Dao et al., 2022).
+To efficiently train the computationally intensive models, we simultaneously employ DeepSpeed training with ZeRO-3 stage [44] and flash attention [6].
 We trained the 7B model on 16 NVIDIA A800 80GB GPUs.
-For the inference stage, we employed the vLLM framework (Kwon et al., 2023) with greedy decoding, setting the temperature to 0.
+For the inference stage, we employed the vLLM framework [23] with greedy decoding, setting the temperature to 0.
 Only one GPU was utilized during the inference phase.
 We only evaluate the instruction-tuned LLaMA model for rating prediction task since it is not applicable for directly making top-$k$ recommendations.
-[…]
-Table 4. Performance achieved by different methods in rating prediction task.
+We implement the models for rating prediction task using the DeepCTR-Torch [^3] https://github.com/shenweichen/DeepCTR-Torch library.
+For the top-$k$ recommendation task, we utilize the SELFRec [^4] https://github.com/Coder-Yu/SELFRec library [65] for implementation.
+As for the hyper-parameter settings, $\alpha_{1}$ and $\alpha_{2}$ are selected from $\{$0.1, 0.3, 0.5, 0.7, 0.9}\} respectively for all experiments.
+$\mathcal{C}$ is fixed to 1 and $k^{\prime}$ is set to 10.
+We repeat the experiment five times and calculate the average.
+We report the best results obtained when the ranking method is selected from pointwise and listwise ranking.
+For all experiments, the best results are highlighted in boldfaces.
+* indicates the statistical significance for $p\leq 0.05$ compared to the best baseline method based on the paired t-test.
+Improv.
+denotes the improvement of our method over the best baseline method.
+IV-B Main Results (RQ1)
+We conducted an extensive evaluation of our proposed Llama4Rec and the baseline methods on three datasets to assess the model’s performance under diverse recommendation scenarios.
+The experiment results for direct recommendation, sequential recommendation, and rating prediction are shown in Table II, Table IV, and Table V, respectively.
+We have the following key observations.
+  • As shown in Table II, Llama4Rec consistently outperforms baseline methods in almost all scenarios, with particularly significant improvements observed in the direct recommendation task.
+Moreover, our findings reveal that direct instruction fine-tuning LLMs for recommendation tasks does not consistently yield promising performance.
+These results highlight the effectiveness of integrating LLMs into conventional recommendation models, underscoring the importance of incorporating the mechanism that utilizes instruction-tuned LLM to mutually augment and adaptively aggregate with conventional recommendation models.
+  • We observe a performance decline when dealing with a larger number of candidates (i.e., $k$) in direct and sequential recommendation scenario, as the improvement for $k=5$ is usually less than for $k=3$.
+This suggests a limitation of the LLM, where ranking and recommending the correct items become more challenging as the candidate pool increases.
+This cause the performance drops in some cases.
+This may be attributed to the “lost-in-the-middle” problem [30], where LLMs tend to focus more on the beginning and ending of sequences while neglecting the middle portion.
+This behavior contributes to a decline in performance as the size of the candidate set increases.
+Besides, larger candidate sets inherently increase the complexity of the task, requiring the model to distribute its attention across more options.
+Smaller or less powerful LLMs may struggle to maintain this balance, leading to reduced performance.
+To address this issue, we propose several possible remedies.
+First, employing more powerful LLMs can help mitigate the problems associated with larger candidate sets, as these models are better equipped to handle longer sequences.
+Second, Splitting large candidate sets into smaller subsets (e.g., partitioning large $k$ into smaller chunks) and aggregating the results can help the model maintain focus and improve accuracy for higher $k$ values.
+  • In the scenario of the rating prediction task in Table V, while the instruction-tuned LLaMA model significantly underperforms when compared to conventional recommendation models, integrating the LLM yields a marked performance improvement.
+This suggests that the LLM and conventional recommendation models learn distinct aspects of information.
+Consequently, integrating the LLM with conventional recommendation models could enhance recommendation performance.
+  • In the context of direct recommendation and sequential recommendation recommendations in Table II and IV , Llama4Rec exhibits a more pronounced improvement for direct recommendations task.
+In addition, a more significant enhancement is observed on the Bookcrossing dataset, which can be attributed to the more fine-grained and distinguishable rating of the Bookcrossing dataset.
+  • Besides, we discuss the computational efficiency.
+Additional training with augmented data is required in the Llama4Rec framework, which presents an extra cost.
+In the current experimental setup, we train a new model from scratch.
+However, this process could be optimized by continuing to train a previously tuned model, thereby reducing time costs.
+Additionally, in our experiment, we observed that training the LLaMA-2 7B model with around 25K instructions on 16 A800 GPUs with 2500 steps took approximately 1.94 hours.
+The inference time for each instruction averaged about 17 instructions per second, translating to a requirement of around 0.059 seconds per item for computation by a single NVIDIA A800 GPU.
+TABLE V: Performance achieved by different methods in rating prediction task.
 Backbone Method ML-100K ML-1M BookCrossing RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ LLaMA IFT 1.2792 0.8940 1.2302 0.8770 2.0152 1.3782 DeepFM Base 1.0487 0.8082 0.9455 0.7409 1.7738 1.3554 Llama4Rec 1.0306* 0.7987* 0.9360* 0.7321* 1.6958* 1.2843* Impro.
@@
 1.25% 0.75% 0.94% 1.05% 6.43% 6.96%
-5.2. Main Results (RQ1)
-We conducted an extensive evaluation of our proposed Llama4Rec and the baseline methods on three datasets to assess the model’s performance under diverse recommendation scenarios.
-The experiment results for rating prediction, direct recommendation, and sequential recommendation are shown in Table 2, Table 3, and Table 4, respectively.
-We have the following key observations.
-  • Llama4Rec consistently outperforms baseline methods in almost all scenarios, with particularly significant improvements observed in the direct recommendation task.
-Moreover, our findings reveal that direct instruction fine-tuning LLMs for recommendation tasks does not consistently yield promising performance.
-These results highlight the effectiveness of integrating LLMs into conventional recommendation models, underscoring the importance of incorporating the mechanism that utilizes instruction-tuned LLM to mutually augment and adaptively aggregate with conventional recommendation models.
-  • In the scenario of the rating prediction task, while the instruction-tuned LLaMA model significantly underperforms when compared to conventional recommendation models, integrating the LLM yields a marked performance improvement.
-This suggests that the LLM and conventional recommendation models learn distinct aspects of information.
-Consequently, integrating the LLM with conventional recommendation models could enhance recommendation performance.
-  • In the context of top-$k$ recommendations, Llama4Rec exhibits a more pronounced improvement for direct recommendations task.
-In addition, a more significant enhancement is observed on the Bookcrossing dataset, which can be attributed to the more fine-grained and distinguishable rating of the Bookcrossing dataset.
-Table 5. Ablation study on key components of Llama4Rec on the ML-1M dataset.
+TABLE VI: Ablation study on key components of Llama4Rec on the ML-1M and Bookcrossing dataset.
 Models ML-1M BookCrossing H@3 $\uparrow$ N@3 $\uparrow$ H@3 $\uparrow$ N@3 $\uparrow$ LightGCN 0.0283 (-) 0.0203 (-) 0.0358 (-) 0.0272 (-) IFT 0.0268 (-5.30%) 0.0193 (-4.93%) 0.0287 (-19.84%) 0.0202 (-25.74%) Llama4Rec w/o DA 0.0294 (+3.89%) 0.0209 (+2.96%) 0.0408 (+13.97%) 0.0319 (+17.28%) Llama4Rec w/o PA 0.0277 (-2.12%) 0.0199 (-1.97%) 0.0372 (+3.92%) 0.0279 (+2.57%) Llama4Rec w/o AA 0.0298 (+5.30%) 0.0218 (+7.39%) 0.0429 (+19.83%) 0.0332 (+22.06%) Llama4Rec 0.0304 (+7.42%) 0.0222 (+9.36%) 0.0434 (+21.23%) 0.0338 (+24.26%)
-5.3. Ablation Study (RQ2)
+IV-C Ablation Study (RQ2)
 We conducted an ablation study to analyze the contributions of different components in our model.
-Table 5 summarizes the results of the ablation studies across three variants on the ML-1M dataset.
+Table VI summarizes the results of the ablation studies across three variants on the ML-1M and Bookcrossing dataset.
 It is evident that the full model performs considerably better than all its variants, indicating that all the main components contribute significantly to overall performance improvement.
-Moreover, compared to the conventional model, the instruction-tuned LLM does not achieve superior results, underscoring the importance of model aggregation.
+Moreover, compared to the conventional model ( e.g., LightGCN ), the instruction-tuned LLM ( e.g., IFT ) does not achieve superior results, underscoring the importance of model aggregation.
 We further analyze the specific impact of each component, and our observations are as follows:
@@
 This indicates the module’s capacity to mitigate data sparsity and long-tail problem, consequently enhancing model performance.
-  • w/o Prompt Augmentation (w/o PA): In this variant, we remove the prompt augmentation component, a crucial element of the proposed framework.
+  • w/o Prompt Augmentation (w/o PA): In this variant, we remove the prompt augmentation component.
 Experimental results demonstrate a significant degradation in model performance when this module is excluded, thereby validating its essential role.
@@
 Experimental results demonstrate a drop in model performance, underscoring the significance of accounting for the user’s long-tail coefficient and employing adaptive aggregation.
-5.4. Hyper-parameter Study (RQ3)
-5.4.1. Analysis of Hyper-parameters $\alpha_{1}$ and $\alpha_{2}$
+IV-D Hyper-parameter Study (RQ3)
 We conducted an analysis of the effects of hyper-parameters $\alpha_{1}$ and $\alpha_{2}$.
@@
 This observation is consistent with the principle of adaptive aggregation, which emphasizes the importance of assigning suitable weights to tail users.
-Figure 3. Impact of hyper-parameters $\alpha_{1}$ and $\alpha_{2}$ on ML-1M dataset with backbone model LightGCN.
-Figure 4. Performance comparison w.r.t different LLaMA-2 size for training Llama4Rec on the Bookcrossing dataset.
-Figure 5. Performance comparison w.r.t different numbers of instructions for training Llama4Rec on the ML-1M dataset.
-5.4.2. Analysis of Model Scaling.
+Figure 4: Impact of hyper-parameters $\alpha_{1}$ and $\alpha_{2}$ on ML-1M dataset with backbone model LightGCN.
+IV-E Model Analysis (RQ4)
+Figure 5: Performance comparison w.r.t different LLaMA-2 size for training Llama4Rec on the Bookcrossing dataset.
+Figure 6: Performance comparison w.r.t different numbers of instructions for training Llama4Rec on the ML-1M dataset.
+IV-E1 Analysis of Model Scaling
 We further instruction-tuned the LLaMA-2 model with different model size.
@@
 It underscores the importance of considering the trade-off between model size, performance gain, and resource efficiency in the design and application of large language models.
-5.4.3. Analysis of Data Scaling.
+IV-E2 Analysis of Data Scaling
 We evaluated the effect of data size on LLM training by varying the number of instructions in the instruction-tuning dataset.
@@
 This highlights the importance of utilizing larger and more diverse datasets for instruction tuning LLMs to optimize performance.
-5.5. Further Discussion
-In this part, we discuss about the computational efficiency and future improvements.
-In the Llama4Rec framework, additional training with augmented data is required, which may present a potential limitation.
-In the current experimental setup, we train a new model from scratch.
-However, this process could be optimized by continuing to train a previously tuned model, thereby reducing time costs.
-Additionally, in our experiment, we observed that training the LLaMA-2 7B model with around 25K instructions on 16 A800 GPUs with 2500 steps took approximately 1.94 hours.
-The inference time for each instruction averaged about 17 instructions per second, translating to a requirement of around 0.059 seconds per item for computation by a single A800 GPU.
+IV-E3 Comparison with the GPT Model
+We compare our Llama4Rec with the GPT model, specifically, the GPT-3.5-turbo [^6] https://platform.openai.com/docs/models/gpt-3-5 model.
+We choose the performance metric hit ratio for comparison.
+A random selection of 100 listwise ranking task samples from the Bookcrossing dataset with backbone model CLSRec served as the benchmark for GPT model.
+The experimental result is shown in Figure 7.
+The experimental result shows Llama4Rec win over the GPT model in 45% of the tasks, tied in 25%, and fell short in 30% of the times.
+This impressive result emphasizes the crucial role of instruction tuning in aligning general-purpose LLMs specifically for recommendation tasks.
+Figure 7: Comparison between our instruction-tuned Llama4Rec with the GPT-3.5-turbo model.
+Tie refers to cases where the proposed framework and GPT achieve identical performance metrics, specifically the hit ratio.
+V Further Discussion
+Model Efficiency.
 This training and inference duration significantly exceeds that of conventional recommendation models, highlighting the limitations of current LLM-based recommender systems.
@@
 This improvement might make the integration of LLMs into large-scale recommender systems feasible, especially by leveraging many GPUs for parallel computation.
-6. Conclusion and future work
+Linear Combination.
+In this work, we chose a heuristic-based linear interpolation method due to its simplicity and effectiveness, as evidenced by our experiments.
+Training a neural network to determine optimal aggregation weights would require substantial labeled data to accurately capture the relationship between input factors and performance, which is resource-intensive and impractical for this study.
+In future work, we plan to explore advanced techniques like genetic programming or reinforcement learning for better modeling of complex factor interactions and enhance nonlinear aggregation effects through dynamic weight optimization.
+Hallucination.
+To ensure the accuracy of attributes generated by the LLM, we focused on well-known movies and books, utilizing their reliable metadata.
+For the evaluation process, we use a two-step approach.
+First, attributes are generated using the LLaMA model.
+Then, we employ another advanced LLM, DeepSeek-V3 [28], as an evaluator to assess the correctness of the generated attributes.
+This approach, referred to as the ”LLM-as-a-Judge” paradigm [73], allows us to systematically evaluate the alignment between generated attributes.
+In cases of disagreement between the evaluator and the generated attributes, we conduct manual evaluations to resolve conflicts, ensuring a higher degree of reliability and accuracy.
+To maintain consistency in our analysis, we randomly sampled 50 items from each dataset and utilized greedy decoding during generation to reduce randomness.
+For ML-100K and ML-1M, we evaluated the generated ”Movie Director” attribute, while for BookCrossing, we evaluated the generated ”Book Genres” attribute.
+The quantitative results of this evaluation process are presented in Figure 8.
+Figure 8: Performance of LLM-generated attributes compared to ground-truth data.
+From Figure 8, we observe that the generated attributes demonstrate a high level of agreement with the ground-truth data, exceeding 80% across all datasets and reaching of 90% in ML-1M.
+However, two common types of errors are observed.
+First, in some cases, the Llama model honestly admits it does not know the answer.
+This behavior could serve as a useful indicator during model training, suggesting that the corresponding movie might not be widely known.
+Second, in other cases, the LLM hallucinates and provides entirely incorrect answers.
+While these errors are undesirable, they may serve as informative noise, potentially contributing to improved model robustness.
+We attribute the high agreement observed to the extensive pretraining of the LLMs on vast amounts of world knowledge, which enables them to generate attributes with a high degree of correctness.
+However, smaller models are more prone to errors or hallucinations, while larger models generally exhibit better performance due to their increased parameter count, broader training data coverage, and reduced hallucination tendencies.
+The world knowledge embedded in LLMs provides a strong foundation for attribute generation.
+However, to further improve accuracy and mitigate the risk of hallucination, we identify several promising future directions.
+These include integrating Retrieval-Augmented Generation (RAG) [25] to ground the model in external, verifiable knowledge sources, leveraging structured Knowledge Graphs [17], or employing next-generation, more powerful foundation models to mitigate these issues.
+VI Conclusion and future work
 In this study, we present Llama4Rec, a general and model-agnostic framework tailored to facilitate mutual augmentation between conventional recommendation models and LLMs through data augmentation and prompt augmentation.
@@
 While our current approach focuses on mutual augmentation within a single step, our future work will explore expanding mutual augmentation in an iterative manner, potentially unlocking further improvements in model performance.
+VII Acknowledgement
+This work was supported in part by the Research Grants Council of the Hong Kong SAR under Grant GRF 11217823, 11216225, and Collaborative Research Fund C1042-23GF, the National Natural Science Foundation of China under Grant 62371411, InnoHK initiative, the Government of the HKSAR, Laboratory for AI-Powered Financial Technologies.